### v1

# Harm Amplification in Text-to-Image Models

Abstract
Warning: The content of this paper as well as some blurred images shown may include references to nudity, sexualization, violence, and gore.
Text-to-image (T2I) models have emerged as a significant advancement in generative AI; however, there exist safety concerns regarding their potential to produce harmful image outputs even when users input seemingly safe prompts. This phenomenon, where T2I models generate harmful representations that were not explicit in the input, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms. Our paper addresses this issue by first introducing a formal definition for this phenomenon, termed harm amplification. We further contribute to the field by developing methodologies to quantify harm amplification in which we consider the harm of the model output in the context of user input. We then empirically examine how to apply these different methodologies to simulate real-world deployment scenarios including a quantification of disparate impacts across genders resulting from harm amplification. Together, our work aims to offer researchers tools to comprehensively address safety challenges in T2I systems and contribute to the responsible deployment of generative AI models.
1 Introduction
Generative text-to-image (T2I) systems allow users to create new image content in response to a text input, learning patterns and relationships from large-scale training data AI (2023); Inc. (2022); AI (2022). As all data are reflections of the social world Gitelman (2013), training data can mirror the harmful stereotypes and social inequalities that exist offline Birhane et al. (2023, 2021); Kirk et al. (2021); Wang et al. (2022), which are subsequently replicated in generated imagery Cho et al. (2023). Responsible AI scholarship illustrates how content from T2I systems can reflect and amplify the harmful representations that exist offline Qu et al. (2023); Qadri et al. (2023); Rando et al. (2022). Harms from generative AI — like all algorithmic systems — occur from the interplay of technical system components and social power dynamics Shelby et al. (2023); Green and Viljoen (2020). Thus, effective harm reduction work requires practitioners to develop sociotechnical safety requirements, conduct evaluations, and implement strategies to proactively minimize harmful representations prior to launch Rismani et al. (2023a); Dobbe (2022); Solaiman et al. (2023).
Responsible AI research has cultivated safety techniques to curb the intentional or unintentional generation of harmful representations, including through reinforcement learning Daniels-Koch and Freedman (2022), dataset remediation Birhane et al. (2023), red teaming Parrish et al. (2023); Rando et al. (2022), model input and output filters Hao et al. (2023), and techniques to suppress unsafe content generation Schramowski et al. (2023). These component safety techniques play a vital role in reducing the likelihood a model will generate harmful representations, but do not aid the empirical measurement of defined harms, which is necessary to assess the effectiveness of safety techniques. Moreover, multimodal systems raise novel evaluation challenges, as harmful representations may result from computational, contextual, and compositional risks between text-image pairings with visual content introducing “additional shock factor[s]" (Weidinger et al., 2023, p. 13). While responsible AI scholarship and practice recognizes the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities" (Shelby et al., 2023, p. 733), limited research has assessed the effectiveness of multimodal safety techniques for a given harm that “could guide the development of evaluation metrics and inform responsible model release" (Saharia et al., 2022, p. 10). A precise definition of harm amplification can strengthen multimodal safety evaluation work enabling deeper understanding of the effectiveness of evaluation techniques.
We address these challenges by formalizing a definition of harm amplification. Further, we offer three statistical methods to examine representational harms that have been defined for a given context, and identify instances of harm amplification in T2I systems. We test this on a large-scale T2I measurement dataset designed to simulate real-world model deployment. In summary, we make the following research contributions:
  • A formal definition of harm amplification for T2I models as occurring when the generated image from a T2I model reflects harmful representations that were not explicit in the text input.
  • Three methods for quantifying and measuring harm amplification instances in T2I models. These methods include using other models such as safety classifiers and CLIP Radford et al. (2021) to determine whether the output image is significantly more harmful than the input prompt.
  • An empirical examination of how to employ the different methodologies to simulate real-world model deployment, which can dynamically inform practitioners’ safety frameworks. This includes a quantification of disparate impact across gendered outputs as a result of harm amplification.
In what follows, we first situate our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification (Section 2); followed by a description of our novel methodologies (Section 4). We then detail our experiments (Section 5) and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems (Section 6), arguing for greater attention to harm amplification in responsible AI research and practice.
2 Related Work
This research lies at the intersection and builds on responsible AI research on harm, safety, and multimodal generative AI. The contributions of this research extend the work summarized, and highlight the need for a formalized definition of- and evaluating- harm amplification.
2.1 Sociotechnical Safety in T2I Systems and Harm Reduction
Concerns about “safety" in AI systems encompass a range of diverse sociotechnical issues that directly or indirectly shape the well-being of people or the environment. As such, safety is contingent and contextual, and an emergent characteristic or property of AI systems Leveson (2020). While approaches to safety in AI systems vary Ahmed et al. (2023), a key concern of a systems approach is to identify system-level hazards Dobbe (2022) and integrate safety considerations into the development process (see: Leveson (2016)) to minimize the likelihood that harms will arise from system use or failure Dobbe (2022). Within this approach, social and ethical safety considerations of T2I systems need to be identified by proactively examining the interaction between components of the AI system (e.g., technical capabilities, potential failures, context of use, intentions of users or other human actors) Rismani et al. (2023a); Weidinger et al. (2023). In this way, safety considerations are reciprocally shaped by social ideas and dynamics that are encoded into AI systems Dobbe (2022); Rismani et al. (2023b). This concept has been described by Ruha Benjamin Benjamin (2019) as the “duplicity of technology," in which AI systems often default to or replicate social power dynamics. Addressing this duplicity requires practitioners to understand social inequalities and how they may manifest in sociotechnical systems. Without intentional consideration and intervention, T2I systems can easily amplify harmful representations Qadri et al. (2023); Bianchi et al. (2023).
2.1.1 AI Governance and Mitigating “Unsafe" Content
An effective governance approach to cultivating “safe" AI systems is to develop clear safety requirements Leveson (2020), with attention to the sociotechnical nature of AI systems Dobbe (2022) and potential harms Shelby et al. (2023). There is increasing recognition of representational harms in T2I systems Cho et al. (2023); Díaz et al. (2023); Bird et al. (2023), which are the socially constructed beliefs about different social groups that reinforce unjust hierarchies Wang et al. (2022); Barocas et al. (2017, 2019), and manifest in different ways (see: Katzman et al. (2021)). Representational harms are sociotechnical in nature Shelby et al. (2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations Birhane et al. (2021). As such, an increasingly common safety requirement (and social and regulatory expectation) Solaiman (2023); Dias Oliva (2020) is reducing the likelihood that T2I systems will amplify harmful representations — such as demeaning gender, racial, disability, or cultural stereotypes Qadri et al. (2023); Bianchi et al. (2023); Ungless et al. (2023); Naik and Nushi (2023); Bennett et al. (2021), particularly when model inputs are benign, unspecified, or do not contain qualifiers (i.e., "irrespective of gender") (Bansal et al., 2022, p. 1359). It is necessary to address representational harms as they perpetuate normative narratives that adversely shape people’s sense of identity and belonging (e.g., Bennett and Keyes (2020); Costanza-Chock (2018); Mengesha et al. (2021)) and algorithmic systems can scale the visibility of harmful representations reinforcing their social power Phillips (2018).
2.1.2 Gender, Social Inequality, and Representation
Gender inequalities are a persistent global problem, which can be exacerbated by AI technologies Lütz (2023). Gender is a social construct through which a society defines expectations around masculinity and femininity, or outside this binary Lorber (1996); West and Zimmerman (1987); West and Fenstermaker (1995). As a social construction, gender ideologies are created and changed to meet social, political, and economic needs Martin (2004). They are an enduring major component structuring individual, interactional, and institutional power relations Risman (2004). A dominant gender construct is the sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment Fredrickson and Roberts (1997); Szymanski et al. (2011). Scholars have long highlighted how constructed ideas about gender difference are used to justify stratification and inequality Lorber (1994), and that gender ideologies intersect with other axes of inequality, including race and ethnicity, socioeconomic status, sexuality, and disability among others. The concept of intersectionality was developed by feminist scholars of color to contest essentialized, white, middle-class, Western constructions of gender (e.g., Anzaldúa (1987); Crenshaw (1991b, a); King (1988); Mohanty (2003)), revealing how gendered representations are informed by intersecting gender, race, and social class ideologies within a “matrix of domination" Collins (1990). Computing scholars have examined gendered stereotypes, tropes, and patterns of over/under representation in algorithms Buolamwini and Gebru (2018), datasets Zhao et al. (2017); Meister et al. (2023); Garrido-Muñoz et al. (2023), and interfaces Kay et al. (2015); Krieg et al. (2023), demonstrating persistent stereotype exaggeration Kay et al. (2015). Gendered stereotypes affect the performance of real world applications, including online advertising Sweeney (2013), toxic language detection Park et al. (2018), machine translation and language technologies Font and Costa-Jussà (2019); Dev et al. (2021); Vanmassenhove et al. (2018), search engines Kopeinik et al. (2023); Kay et al. (2015); Albawardi and Jones (2023), and T2I systems Bianchi et al. (2023); Naik and Nushi (2023).
2.2 Stereotypes & Amplification of Societal Biases in T2I Systems
The representational harm phenomenon in T2I models has been clearly identified (e.g., Hao et al. (2023); Bird et al. (2023); Bianchi et al. (2023); Cho et al. (2023); Ungless et al. (2023); Wolfe et al. (2023); Yu et al. (2022); Jain et al. (2021); Choi et al. (2020)). Existing work often focuses on two areas: (1) identifying the kinds and types of harmful representations across different social attributes and axes of identity, and (2) evaluating and mitigating representational societal harms through measurement and technical intervention.
2.2.1 Amplifying Harmful Societal Representations
Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone Jain et al. (2021); Choi et al. (2020). Jain et al. Jain et al. (2021) demonstrate how generated images have a “distributional shift" away from real world data, in which generative AI systems exacerbate extant social inequalities and patterns of over/under-representation. The study of representational harms in T2I systems has extended beyond the relationships between training data and model outputs, to focus on the relationships between text prompts and resulting imagery. Bianchi and colleagues Bianchi et al. (2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts. They show using Stable Diffusion, for example, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p. 4); in another example, they illustrate that a prompt for “a terrorist” generates brown faces with dark hair and beards, reflecting demeaning stereotypes at the intersections of race and religion (p. 3). Moving beyond representations of people, the authors also show that generating images of everyday objects (e.g., “a house") reflects Western and middle-class stereotypes (p. 3). Adding specific targeted identity terms (i.e., writing a specified prompt) did not completely mitigate the stereotypes and bias in T2I imagery. Qadri et al. (Qadri et al., 2023, p. 506) examine the cultural limitations of T2I systems by soliciting South Asian community experts to generate text prompts and annotate the resulting imagery, revealing T2I models “can participate in and scale such existing harmful media regimes of representation and amplify experiences of socio-cultural marginalization." Similarly, Naik et al. Naik and Nushi (2023) explore exacerbation of social inequalities along axes of (perceived) gender, age, race, and geographical location, showing state-of-the-art T2I systems amplify real world inequities when compared to U.S. Bureau of Labor Statistics. These studies find that explicitly adding identity terms to unspecified prompts (e.g., “female security guard" Naik and Nushi (2023)) does not effectively mitigate the presence of stereotypes in generated imagery Bianchi et al. (2023).
2.2.2 Evaluating Representational Societal Biases
A parallel body of research focuses not just on identifying the presence and kinds of harmful stereotypes that T2I systems can generate, but also evaluating them. Luccioni and colleagues Luccioni et al. (2023) examine whether representational societal harms in each modality amplify or compound one another by evaluating images generated from two synthetic text prompt datasets: one with gender and ethnicity terms and one with profession identifiers. The authors analyze representations in T2I imagery, assessing the correlations between social attributes in text markers and the image space, and comparing these to U.S. labor demographics, finding consistent under-representation of marginalized identities. Wang et al. Wang et al. (2023) employ the Text-to-Image Association Test to measure and quantify the presence of societal bias in T2I models while measuring stereotype amplification in occupations across genders. Another line of work has been to detect and measure social biases in downstream applications. Perera et al. Perera and Patel (2023) explore social bias with respect to gender, race, and age in the diffusion-based face generation models, finding that diffusion models exacerbate patterns of over- and under-representation, while Saravanan et al. (2023) conducts an intersectional analysis of racialized gender stereotypes and professions to support evaluation of T2I models. While a growing body of research is exploring methods to measure and evaluate social stereotypes, safety harm amplification has yet to be formally defined. Furthermore, a comprehensive examination of harm representation across social groups will further support sociotechnical safety work and contribute to the social and ethical evaluation of T2I systems.
3 Defining Harm Amplification
Our paper focuses on examining representational harms, referring to the socially constructed beliefs about different social groups that reinforce unjust hierarchies Wang et al. (2022); Barocas et al. (2017); Katzman et al. (2021); Barocas et al. (2019), in a specific set of generative AI: T2I systems. First, we define “unsafe" content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements. What could be considered “unsafe" or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations Rismani et al. (2023a); Dobbe (2022); Leveson (2020). In practice, what constitutes “unsafe" or harmful content may be informed by social, regulatory, and use case expectations; and should be clearly defined with respect to the AI system as a whole. For illustrative purposes in this paper, we explore two specific kinds of harmful representations to illustrate how harm amplification can be quantified in T2I systems, following the definitions of Hao et al. Hao et al. (2023):
  • Sexually explicit content: Depictions of explicit or graphic sexual acts, nudity (beyond human anatomy and presented in a sexually suggestive manner), sexualized body parts, or sex toys.
  • Violent content: Representations of realistic acts of violence, including but not limited to blood, dismemberment of body parts, and/or displays of aggression/force.
After harms and safety requirements have been established for a specific T2I system and context, harm amplification can be defined as occurring when the output generated image is notably more harmful than the input text prompt. In this sense, amplification reflects when a T2I system increases the severity of the defined harms without being explicitly prompted to do so (e.g., sexually objectifying women in generated imagery). Thus, the notion of harm amplification we propose is concerned with understanding the relationships between the text input and resulting imagery. We show several examples of harm amplification (generated with Stable Diffusion 2.1 Rombach et al. (2022)) in Figure 1 for sexually explicit (Fig 1A) and violent content (Fig 1B) identified by our measurement techniques.
Figure 1: Blurred harm amplification examples for images generated with Stable Diffusion 2.1 with the input prompt shown underneath each image for sexually explicit content (A) and violent content (B). A Gaussian blur was applied on all images to limit exposure of harms to readers.
Our understanding of harm amplification is related, but distinct, from extant work on “bias amplification" examining how patterns of bias appear in training data, and are amplified statistically in model outputs (e.g., Zhao et al. (2017); Wang and Russakovsky (2021); Zhao et al. (2023); Lloyd (2018); Taori and Hashimoto (2023); Mehrabi et al. (2021)). In this literature, “bias amplification" reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics" (Hall et al., 2022a, p. 1), often focusing on binary classification models (e.g., Leino et al. (2019); Wang et al. (2022); Hall et al. (2022b)) and correlations Zhao et al. (2017). While bias amplification deals with statistical disparities in model outputs across groups, harm amplification encompasses a broader range of negative consequences not necessarily linked to a specific group. Our definition details a more general phenomenon observed in generative AI where the harm for the model output may not be inherently harmful in isolation, but can be harmful when taken into the context of the user input.
4 Methods to Measure Harm Amplification
Quantifying harm amplification is a crucial component in assessing “safety" in T2I systems with respect to defined unsafe content. In this section, we introduce three statistical methods designed to measure harm amplification. Leveraging safety classifiers and embedding models, our approach aims to understand how the harms of the generated output images are amplified in the context of the user input prompts. Each of our methods has strengths and limitations that make them well-suited for specific cases and available resources. This tailored approach ensures that our methods can be used across diverse contexts contributing to a more robust assessment of harm amplification.
4.1 Method 1: Distribution-Based Thresholds for Discretized Text Harm Scores
The Distribution-Based Thresholds method utilizes (1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and (2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image). Consider a set of text harm scores ($T$) and image harm scores ($I$). We first discretize $T$ into $n$ even buckets ($B_{0},B_{1},\text{…},B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
$B_{j}=\{t\in T:l_{j}<t\leq l_{j+1}\},\quad j=0,1,...,n$ (1)
For each $B_{j}$, the corresponding image harm scores ($I_{j}$) form a distribution ($D_{j}$). Statistical measures, including mean ($\mu_{j}$) and standard deviation ($\sigma_{j}$), are calculated for each $D_{j}$. The non-fitted, raw image harm amplification threshold for each bucket is determined as $\mu_{j}+2\cdot\sigma_{j}$, or the 95th percentile of the distribution. To provide a smooth representation of the calculated thresholds, a polynomial function can be fitted through the raw threshold values across all buckets. In our experiments, we fit a first-degree polynomial with the corresponding equation to determine the image harm amplification threshold $HarmAmpThresh_{j}$ for each text harm bucket $B_{j}$:
$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0}$ (2)
where $\text{where }y_{j}=\mu_{j}+2\cdot\sigma_{j}$
For example, consider a T2I prompt-image dataset where we have image harm scores $I$ and text harm scores $T$. Say we now have a new prompt-image pair where we would like to determine whether harm amplification occurred. Given an image harm score (i.e., $i=0.5$) and text harm score (i.e., $t=0.15$), we categorize the text harm score into a specific bucket (i.e., $B_{1}=0.1<t\leq 0.2$). For that given bucket, the image harm amplification threshold (i.e., $HarmAmpThresh_{1}$) is obtained by fitting a 1-degree polynomial to the 95th percentile threshold for a prior image harm distribution in each bucket. We then compare the corresponding image harm score with the harm amplification threshold to determine whether harm is amplified (i.e., $i>HarmAmpThresh_{1}$).
This method of Distribution-Based Thresholds is best suited to situations where text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier). Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
4.2 Method 2: Bucket Flip for Discretized Text and Image Scores
Utilizing text and image harm safety scores similar to Method 1 (Section 4.1), the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text. Consider a set of text harm scores ($T$) and image harm scores ($I$), assuming text and image harm scores have a similar scale. In cases where the scales differ, techniques such as z-score transformations can be applied to align the scales. We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\text{…},B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$. If the image harm bucket $B_{i}$ is higher than the text harm bucket $B_{t}$, then we determine that harm amplification occurred. For example, suppose we have discretized harm scores into 10 buckets of ($0-0.1,0.1-0.2,\text{…},0.9-1.0$). Given an image harm score (i.e., $i=0.45$) and text harm score (i.e., $t=0.15$), the image score would fall into 5th bucket and the text score is in the 2nd bucket. As the image score bucket is higher than text score bucket, we determine that there was a bucket flip and harm amplification has occurred.
The Bucket Flip method is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores. In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical approach for identifying harm amplification.
4.3 Method 3: Image-Text Co-embedding Based Harm Scores
While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many people may not have the resources to label such data and thus may not be able to train the necessary safety classifiers. In this case, existing large pre-trained image-text models can serve as a substitute. Large pre-trained image-text models, such as CLIP Radford et al. (2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space. Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s embedding, $i$, is to a harmful concept in the embedding space than the prompt’s embedding, $p$, is. Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the harm sexually explicitly, we can use the words “sexualized", “porn", etc.), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$. These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space. Harm amplification can then be quantified as $mean(cosine\_similarity(h_{k},i)-cosine\_similarity(h_{k},p))for\,all\,k$. This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt.
5 Experiments
In this section, we demonstrate how to employ the methods that we outlined in Section 4 on a large and varied T2I measurement dataset designed to simulate usage in real-world model deployment to develop a harm amplification criteria for each method. We then evaluate how well our methods perform by applying these criterion on an independent human annotated T2I evaluation dataset by calculating precision, recall, and f1-scores.
5.1 Measurement Dataset & Models
Our measurement dataset consists of 497,157 prompts representing various demographics and prompt categories. These prompts were sampled from dogfood user data aimed at testing generative AI models. By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment. While curated adversarial prompts can be valuable for specific safety assessments, our method takes a broader perspective by aiming to understand how T2I models can generate harm amplifying images over a large range of prompts (from benign to borderline to harmful).
We used Stable Diffusion 2.1 Rombach et al. (2022) to generate images for our prompt measurement dataset. Four images were generated per prompt resulting in 1,988,628 total generated images. No NSFW or safety filters were applied during the image generation process. Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content. Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1. Fairness machine learning classifiers that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1. For the pre-trained image-text model, we use CLIP Radford et al. (2021) and generated embeddings for images and their corresponding prompts.
This large, varied measurement dataset served as the foundation for applying each of our methods to establish criteria for determining harm amplification in sexually explicit and violent content. Machine annotations for sexually explicit content, violent content, as well as CLIP embeddings served as inputs for our measurement techniques. After applying each of our measurement techniques to our measurement dataset, a criteria was obtained for each method:
  1. 1.
    Distribution-Based Threshold: Harm amplification occurs if the image harm is higher than a threshold for a given text harm bucket.
  2. 2.
    Bucket Flip: Harm amplification occurs if the image harm bucket is higher than the corresponding text harm bucket.
  3. 3.
    Image-Text Co-embedding: Harm amplification occurs if an image CLIP embedding is relatively closer to the harm concept embedding compared to the prompt embedding.
5.2 Evaluation Data & Metrics
To create the evaluation dataset, we sourced 742 prompt, image pairs from the Adversarial Nibbler data challenge Alicia Parish (2023). This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 AI (2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) AI (2022). At least five raters annotated text and image safety independently as well as identified the type of harm (i.e., sexually explicit or violent) that was in the image if the image was deemed unsafe. Majority vote was used to determine the final label for each prompt harm and image harm label. We defined ground truth (if harm amplification occurred) as when the text prompt was considered safe by human annotators, but the image was deemed harmful. Ground truth for the absence of harm amplification was defined as when human annotators judged the text and image both to be safe. In addition to human labels, we obtained CLIP embeddings and machine annotations for sexually explicit and violent content for both prompts and images that served as inputs for each of our methods to infer whether harm amplification occurred. Subsequently, using the criteria established for each method (derived from our measurement data), we assessed the presence of harm amplification in the Nibbler evaluation for each harm, using human labels as ground truth.
We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset. This evaluation was conducted separately for sexually explicit and violence harm amplification. Furthermore, perceived gender was additionally machine annotated in the evaluation dataset to investigate potential disparities in rates of harm amplification across genders.
5.3 Experimental Results
5.3.1 Sexually Explicit & Violent Content Results
To quantify harm amplification using the Distribution-Based Threshold method (Section 4.1), we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence. Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix A, Figure 2A). The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket. A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix A, Figure 2B) for each harm type. For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution. Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset. Notably, we observed high precision (0.925) and recall (0.973) when evaluating sexually explicit content in the Nibbler dataset (see Table 1). However, we noted slightly lower precision (0.788) and recall (0.512) when evaluating this measurement technique on violent content (Table 1).
Using the Bucket Flip method (Section 4.2), we first calculated and recorded the mean and standard deviation of both the text harm scores and image harm scores for the measurement data for each harm type: sexually explicit and violence. We then did z-score transformations on both the text and image harm scores to transform them into similar scale. We then subsequently divided that range into 10 even buckets. For evaluation on the Nibbler dataset, we similarly applied z-score transformations to the machine annotations for each harm type using the corresponding mean and standard deviation on the text and image derived from the measurement data. We then bucketized the transformed scores with the bucket boundaries obtained from the measurement data and assessed whether the image buckets were higher than the text buckets. We observed high precision (0.896) and recall (0.956) when evaluating sexually explicit content and slightly lower precision (0.870) and recall (0.488) when evaluating on violent content.
While we expect the Image-Text Co-embedding method (Section 4.3) to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources. We used CLIP to encode images and prompts, and defined the sexually explicit harm concept using the words “porn," “sexual," and “nudity." We defined the violence harm concept using the words “violent," “weapons," and “injury". Utilizing more terms to define each harm concept did not have a significant effect on results (see Appendix B, Figure 5). Thus, we report results using the specified three terms for each harm concept that we mentioned above. Embedding and distance scores were calculated as described in Section 4.3 separately for the measurement data and evaluation data. Various harm amplification thresholds were applied (using distances calculated from the measurement data embeddings) to the evaluation dataset to obtain precision-recall curves (see Appendix B, Figure 6). Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case. For these results we set an operating recall of 0.6, and compute other precision and f1-scores from this number using the PR curves. As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.821, recall of 0.600) than violence harm amplification (precision 0.390, recall 0.600).
Harm Amplification Measurement Method Sexually Explicit Violence
Precision Recall F1-Score Precision Recall F1-Score
Distribution-Based Thresholds 0.925 0.973 0.948 0.788 0.512 0.621
Bucket Flip 0.896 0.956 0.925 0.870 0.488 0.625
Image-Text Co-embedding 0.821 0.600 0.693 0.390 0.600 0.473
Table 1: Harm Amplification measurement methods evaluation on Nibbler data. Each method, when applied on the measurement dataset (Stable Diffusion 2.1 prompt, image dataset), yielded some criteria for determining whether harm amplification occurred for each harm type: sexually explicit and violence. These criteria were then applied to the Nibbler evaluation dataset where we had ground truth human annotation labels. Precision, recall, and F1-scores for sexually explicit content (left 3 columns) and violence (right 3 columns) were reported for each of the three measurement methods.
5.3.2 Gender Fairness
As described in Related Work (Sections 2.1.2 & 2.2.1), the amplification of social stereotypes disproportionately affects marginalized communities Bianchi et al. (2023); Naik and Nushi (2023). We present quantitative measurement of harm amplification — particularly the gendered sexual objectification stereotype Fredrickson and Roberts (1997); Szymanski et al. (2011) — and show that it disproportionately affects images depicting perceived females. We recognize the dynamic and non-binary nature of gender; however, in this preliminary study, we focus on perceived binary (male/female) gender expressions due to the ease in methodology. We also acknowledge that we that cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
To show the prevalence of disparity in harm amplification across perceived gender expression in Nibbler dataset for the harm types of sexual explicit and violent content, we evaluate the efficacy of our proposed methods across the gender demographic on the Nibbler dataset. From the Nibbler dataset, for all images, both the number of people and their perceived gender expression were determined by machine classification Baruah et al. (2022). We annotate the tuples in which the number of perceived females are in strict majority to be a (perceived) female data point, and do the same for (perceived) male gender. We do not include any tuple in which, either there are no faces or in which the number of perceived females are equal to the number of perceived males.
Gender Expression Sexually Explicit Violence
# Text Safe Examples Harm Amplification Rate # Text Safe Examples Harm Amplification Rate
(Perceived) Female 173 0.80 60 0.43
(Perceived) Male 145 0.46 120 0.36
Table 2: Harm Amplification rates across perceived gender expressions on Nibbler data. Significance testing was done using z-test. Statistically significant numbers are in bold ($p<0.01$).
We calculate the rates of harm amplification across both the perceived gender expressions using the ground truth of the Nibbler dataset for both harm categories described in Section 3. Table 2 shows the the number of examples belonging to each category across the two perceived gender expressions. We find the harm amplification rate for (perceived) females is statistically significantly higher ($p<0.01$) for the category of sexually explicit harm amplification. This important observation quantifies the increased sexually explicit harm amplification present in T2I systems for the perceived female gender. Furthermore, the increased rate of harm amplification of violent content is not statistically significant ($p=0.181$).
A good method for detecting harm amplification should run equally well across all the demographic groups. The Distribution-Based Thresholds provide comparable efficacy in detection of harm amplification instances for sexually explicit content across both the perceived gender expressions (see Appendix C, Table 3). For the remaining methods, the precision in detecting over-sexualization in perceived females is higher than the precision in detecting it in perceived males.
6 Discussion & Conclusion
In this paper, we provide a formal definition of harm amplification in T2I models as occurring when the generated output image reflects more harmful representations that were not explicit in the text input (Section 3). Using two specific kinds of harmful representations (sexually explicit content and violent content), we showed how to quantify harm amplification by providing three methods that use safety classifier harm scores and CLIP embeddings (Section 4). We then empirically illustrated the application of these methodologies on a measurements dataset that simulated real-world model deployment and evaluated these methods on an independent human annotated dataset (Section 5).
6.1 Using Data-Driven Methods to Quantify Harm Amplification
Each of our three data-driven methods for quantifying harm amplification offers its own strengths and limitations. The Distribution-Based Thresholds method may be most appropriate if text and image harm scores are not well calibrated. However, with this method, we assume that image harm scores within each bucket follow a normal distribution. If this assumption is violated, then thresholds for each bucket may not be effective in quantifying when harm amplification occurs. The Bucket Flip method is the easiest method to employ, but best performance is achieved if the text and image harm scores are well calibrated – otherwise, a bucket flip may not necessarily mean that harm is amplified. Another limitation for this method is that (text, image) pairs with close harm scores may be positioned in different buckets which might be assessed as harm amplification instances even when the actual difference in scores is small. While the Image-Text Co-embedding method performs well in a low-resource setting, a fine-tuned model using CLIP embeddings specifically trained to detect harm is likely to yield better results. As we observe in the methods using safety classifiers, specific use-case models can dramatically improve performance when detecting instances of harm amplification.
Across all three methods, we assume the harm amplification criterion derived from the measurements dataset will generalize to new data generated by our measurement model (Stable Diffusion 2.1) as well as new data generated by other T2I models. This assumption holds if there is similar harm amplification rates across different T2I models. Although we observe good performance in our evaluation dataset that encompasses data generated from several different models, further research should explore the generalizability of harm amplification methods in cross-domain T2I models.
6.2 Reduced Reliance on Human Annotation
Our novel framework significantly reduces the reliance on human annotations, which may be especially helpful in low-resource settings. By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets. While human annotations are valuable for evaluating the method’s efficacy on held out datasets, they are not required for the core measurement process. Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive. Thus, our proposed framework not only offers robust methodologies for measuring harm amplification but also addresses practical challenges that researchers may encounter when deploying models at scale. Of course, machine annotations — akin to human annotations Davani et al. (2022); Denton et al. (2021) — can also reflect social biases. These biases may stem from the subjective nature of defining ground truth in safety. Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
6.3 Gender Disparities in Harm Amplification
In analyzing rates of harm amplification in the evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content. Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images. These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender. Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts in different sociodemographic groups. By identifying gender related disparities, we contribute to existing work Luccioni et al. (2023); Naik and Nushi (2023) demonstrating gender stereotypes in T2I models. The durability of gender stereotypes in AI systems Buolamwini and Gebru (2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes. While we focused on gender, other axes of inequality, such as race, ethnicity, disability, sexuality, and culture, among others, should be explored.
6.4 Future Work
While our work lays a foundation to better understand and quantify harm amplification in T2I systems, further research is needed to more holistically understand the phenomenon and develop strategies for its reduction in T2I models. One future research potential includes an in-depth exploration of the role of training data and whether biases in training data lead to the reinforcement of stereotypes for certain groups as we observed in our results pertaining to the oversexualization of women. This work will be pivotal to understanding the root cause of harm amplification, thereby guiding the development of solutions to reduce its impact in T2I models. Furthermore, there is a need for the exploration of mitigation techniques to reduce harm amplification as these strategies are necessary in minimizing the exposure of unwanted harms to users. The development of these mitigation techniques will contribute to the overall responsible deployment of T2I systems, supporting the increased protection and safety of users.
7 Acknowledgements
We thank Kathy Meier-Hellstern, Alexander Vorontsov, and Vinodkumar Prabhakaran for their feedback and suggestions. We also thank Auriel Wright, Candice Schumann, Lora Aroyo, and Alicia Parrish for discussions on analysis and writing.
Appendix A Experimental Methods Applied on the Measurement Dataset
Figure 2: Method 1: Distribution-based thresholds applied on the measurements dataset. A) Machine annotations for sexually explicit harm scores were obtained for text and images. Text sexually explicit scores were bucketed into 5 buckets. Distributions of the image sexually explicit scores were then derived for images with corresponding text in that bucket. 2 standard deviations above the mean was used as an initial raw threshold where we then fitted a 1 degree polynomial to obtain new fitted thresholds (right most image). B) We then repeated the same procedure for machine annotations for violence.
Appendix B Evaluation Results on Nibbler Data
Figure 3: Confusion Matrix for Method 1: Distribution-based thresholds on the Nibbler evaluation dataset. Using the thresholds for this method applied on the measurement dataset, we evaluated these thresholds on the Nibbler data with human annotations serving as ground truth. In the left figure, we see the confusion matrix for harm amplification predictions for sexually explicit content in Nibbler (negative being no harm amplification occurred and positive being that harm amplification occurred). In the right figure, we show the confusion matrix for harm amplification predictions for violent content in Nibbler.
Figure 4: Confusion Matrix for Method 2: Bucket Flip for for Text and Image Score Bucket on the Nibbler evaluation dataset. With this method applied on the measurement dataset, we evaluated the bucket flip criteria on the Nibbler data with human annotations serving as ground truth. The left figure shows the confusion matrix for sexually explicit content and the right figure shows the confusion matrix for violent content
Figure 5: Comparing the PR curves with different amount of terms in Method 3. For sexually explicit content, we show no notable difference between the PR curves for 3 terms ("porn", "sexual", "nudity") and 10 terms ("porn", "sexual", "nudity", "erotic", "sex", "breasts", "penis", "vagina", "erotica", and "indecency"). AUC difference was also minimal (3 terms: 0.846, 10 terms: 0.848). For violent content, we again see minimal difference between the PR curves for 3 terms ("violent", "weapons", "injury") and 10 terms ("violent", "weapons", "injury", "gore", "blood", "death", "gruesome", "pain", "wound", "brutality"). Again, AUC difference between 3 terms (0.412) and 10 terms (0.397) was small.
Figure 6: Precision-recall curves for the Image-Text Co-embedding method for sexually explicit content (left) and violent content (right) for the final 3 terms used.
Appendix C Gender Fairness Analysis
Table 3 shows the efficacy of our proposed methods across both the perceived gender expressions for the harm amplification of sexually explicit content. Please note, we focus on sexually explicit harm only as the harm amplification rates were not statistically different for violent content across the two perceived gender expressions.
Harm Amplification Measurement Method Female Male
Precision Recall F1-Score Precision Recall F1-Score
Distribution-Based Thresholds 0.978 0.971 0.974 0.970 0.970 0.970
Bucket Flip 0.985 0.964 0.974 0.753 0.865 0.805
Image-Text Co-embedding 0.978 0.619 0.758 0.760 0.567 0.650
Table 3: Efficacy of the proposed methods across the two perceived gender expressions for sexually explicit harm amplification.

### v2

@@
 Abstract
-Warning: The content of this paper as well as some blurred images shown may include references to nudity, sexualization, violence, and gore.
+Warning: The content of this paper as well as some blurred images shown include references to nudity, sexualization, violence, and gore.
 Text-to-image (T2I) models have emerged as a significant advancement in generative AI; however, there exist safety concerns regarding their potential to produce harmful image outputs even when users input seemingly safe prompts.
 This phenomenon, where T2I models generate harmful representations that were not explicit in the input, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms.
-Our paper addresses this issue by first introducing a formal definition for this phenomenon, termed harm amplification.
-We further contribute to the field by developing methodologies to quantify harm amplification in which we consider the harm of the model output in the context of user input.
+Our paper addresses this issue by formalizing a definition for this phenomenon which we term harm amplification.
+We further contribute to the field by developing a framework of methodologies to quantify harm amplification in which we consider the harm of the model output in the context of user input.
 We then empirically examine how to apply these different methodologies to simulate real-world deployment scenarios including a quantification of disparate impacts across genders resulting from harm amplification.
 Together, our work aims to offer researchers tools to comprehensively address safety challenges in T2I systems and contribute to the responsible deployment of generative AI models.
-1 Introduction
-Generative text-to-image (T2I) systems allow users to create new image content in response to a text input, learning patterns and relationships from large-scale training data AI (2023); Inc.
-(2022); AI (2022).
-As all data are reflections of the social world Gitelman (2013), training data can mirror the harmful stereotypes and social inequalities that exist offline Birhane et al.
-(2023, 2021); Kirk et al.
-(2021); Wang et al.
-(2022), which are subsequently replicated in generated imagery Cho et al.
-(2023).
-Responsible AI scholarship illustrates how content from T2I systems can reflect and amplify the harmful representations that exist offline Qu et al.
-(2023); Qadri et al.
-(2023); Rando et al.
-(2022).
-Harms from generative AI — like all algorithmic systems — occur from the interplay of technical system components and social power dynamics Shelby et al.
-(2023); Green and Viljoen (2020).
-Thus, effective harm reduction work requires practitioners to develop sociotechnical safety requirements, conduct evaluations, and implement strategies to proactively minimize harmful representations prior to launch Rismani et al.
-(2023a); Dobbe (2022); Solaiman et al.
-(2023).
-Responsible AI research has cultivated safety techniques to curb the intentional or unintentional generation of harmful representations, including through reinforcement learning Daniels-Koch and Freedman (2022), dataset remediation Birhane et al.
-(2023), red teaming Parrish et al.
-(2023); Rando et al.
-(2022), model input and output filters Hao et al.
-(2023), and techniques to suppress unsafe content generation Schramowski et al.
-(2023).
-These component safety techniques play a vital role in reducing the likelihood a model will generate harmful representations, but do not aid the empirical measurement of defined harms, which is necessary to assess the effectiveness of safety techniques.
-Moreover, multimodal systems raise novel evaluation challenges, as harmful representations may result from computational, contextual, and compositional risks between text-image pairings with visual content introducing “additional shock factor[s]" (Weidinger et al., 2023, p.
+Introduction
+Generative text-to-image (T2I) systems allow users to create new image content in response to a text input, learning patterns and relationships from large-scale training data (Open AI 2023; Midjourney Inc.
+2022; Stability AI 2022).
+As all data are reflections of the social world (Gitelman 2013), training data can mirror the harmful stereotypes and social inequalities that exist offline (Birhane et al.
+2023; Birhane, Prabhu, and Kahembwe 2021; Kirk et al.
+2021; Wang et al.
+2022), which are subsequently replicated in generated imagery (Cho, Zala, and Bansal 2023).
+Responsible AI scholarship illustrates how content from T2I systems can reflect and amplify the harmful representations or stereotypes of groups that exist offline (Qu et al.
+2023; Qadri et al.
+2023; Rando et al.
+2022b).
+Harms from generative AI — like all algorithmic systems — occur from the interplay of technical system components and social power dynamics (Shelby et al.
+2023; Green and Viljoen 2020).
+Thus, effective harm reduction work requires practitioners to develop sociotechnical safety requirements, conduct evaluations, and implement strategies to proactively minimize harmful representations prior to launch (Rismani et al.
+2023a; Dobbe 2022; Solaiman et al.
+2023).
+Responsible AI research has cultivated safety techniques to curb the generation of unsafe content including through reinforcement learning (Daniels-Koch and Freedman 2022), dataset remediation (Birhane et al.
+2023), red teaming (Parrish et al.
+2023; Rando et al.
+2022b), model input and output filters (Hao et al.
+2023), and techniques to suppress harmful content generation (Schramowski et al.
+2023).
+These component safety techniques play a vital role in reducing the likelihood a model will generate harmful representations.
+However, these techniques do now eliminate all potential harms such as the exposure of unintentional contextualized harms in generative AI systems.
+Consider a scenario where a user requests an image of “people at a party” from a multimodal AI model, and images of women in bikinis dancing are outputted.
+The input prompt and output image in isolation may be considered safe, but when taken together in context of each other, they can expose unintentional harm (oversexualization of women) to the user.
+This scenario underscores the complex challenges posed by multimodal systems.
+Harmful representations may result from computational, contextual, and compositional risks between text-image pairings with visual content introducing “additional shock factor[s]” (Weidinger et al.
+2023, p.
 13).
-While responsible AI scholarship and practice recognizes the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities" (Shelby et al., 2023, p.
-733), limited research has assessed the effectiveness of multimodal safety techniques for a given harm that “could guide the development of evaluation metrics and inform responsible model release" (Saharia et al., 2022, p.
-10).
-A precise definition of harm amplification can strengthen multimodal safety evaluation work enabling deeper understanding of the effectiveness of evaluation techniques.
-We address these challenges by formalizing a definition of harm amplification.
-Further, we offer three statistical methods to examine representational harms that have been defined for a given context, and identify instances of harm amplification in T2I systems.
-We test this on a large-scale T2I measurement dataset designed to simulate real-world model deployment.
+While responsible AI scholarship acknowledges the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities” (Shelby et al.
+2023, p.
+733), a precise definition of multimodal safety harm amplification has yet to be introduced, to our knowledge.
+To address these challenges, we formalize the concept of harm amplification in T2I systems and present a framework to quantify harm amplification.
+This framework employs three statistical methods that can be used in different contexts, including scenarios with: 1) large amounts of data and access to safety classifiers 2) limited data, but having access to safety classifiers, and 3) low resource settings with access only to open source models.
 In summary, we make the following research contributions:
-  • A formal definition of harm amplification for T2I models as occurring when the generated image from a T2I model reflects harmful representations that were not explicit in the text input.
-  • Three methods for quantifying and measuring harm amplification instances in T2I models.
-These methods include using other models such as safety classifiers and CLIP Radford et al.
-(2021) to determine whether the output image is significantly more harmful than the input prompt.
-  • An empirical examination of how to employ the different methodologies to simulate real-world model deployment, which can dynamically inform practitioners’ safety frameworks.
+  • A formal definition of harm amplification for T2I models as occurring when the generated image from a T2I model reflects harmful or unsafe representations that were not explicit in the text input.
+  • A framework that includes three methods for quantifying and measuring harm amplification instances in T2I models.
+These methods include using other pretrained models such as safety classifiers and CLIP (Radford et al.
+2021) to determine whether harm amplification has occurred (i.e., when the output image is more harmful than the user input prompt).
+  • An empirical examination of how to employ each methodology to simulate different real-world model deployment scenarios, which can dynamically inform practitioners’ safety frameworks.
 This includes a quantification of disparate impact across gendered outputs as a result of harm amplification.
-In what follows, we first situate our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification (Section 2); followed by a description of our novel methodologies (Section 4).
-We then detail our experiments (Section 5) and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems (Section 6), arguing for greater attention to harm amplification in responsible AI research and practice.
-2 Related Work
+In what follows, we first orient our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification; followed by a description of our novel methodologies.
+We then detail our experiments and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems, arguing for greater attention to harm amplification in responsible AI research and practice.
+Related Work
 This research lies at the intersection and builds on responsible AI research on harm, safety, and multimodal generative AI.
 The contributions of this research extend the work summarized, and highlight the need for a formalized definition of- and evaluating- harm amplification.
-2.1 Sociotechnical Safety in T2I Systems and Harm Reduction
-Concerns about “safety" in AI systems encompass a range of diverse sociotechnical issues that directly or indirectly shape the well-being of people or the environment.
-As such, safety is contingent and contextual, and an emergent characteristic or property of AI systems Leveson (2020).
-While approaches to safety in AI systems vary Ahmed et al.
-(2023), a key concern of a systems approach is to identify system-level hazards Dobbe (2022) and integrate safety considerations into the development process (see: Leveson (2016)) to minimize the likelihood that harms will arise from system use or failure Dobbe (2022).
-Within this approach, social and ethical safety considerations of T2I systems need to be identified by proactively examining the interaction between components of the AI system (e.g., technical capabilities, potential failures, context of use, intentions of users or other human actors) Rismani et al.
-(2023a); Weidinger et al.
-(2023).
-In this way, safety considerations are reciprocally shaped by social ideas and dynamics that are encoded into AI systems Dobbe (2022); Rismani et al.
-(2023b).
-This concept has been described by Ruha Benjamin Benjamin (2019) as the “duplicity of technology," in which AI systems often default to or replicate social power dynamics.
+Sociotechnical Safety in T2I Systems and Harm Reduction
+Concerns about safety in AI systems encompass a range of diverse sociotechnical issues that directly or indirectly shape the well-being of people or the environment.
+As such, safety is contingent and contextual, and an emergent characteristic or property of AI systems (Leveson 2020).
+While approaches to safety in AI systems vary (Ahmed et al.
+2023), a key concern of a systems approach is to identify system-level hazards (Dobbe 2022) and integrate safety considerations into the development process (Leveson 2016) to minimize the likelihood that harms will arise from system use or failure (Dobbe 2022).
+Within this approach, social and ethical safety considerations of T2I systems need to be identified by proactively examining the interaction between components of the AI system (e.g., technical capabilities, potential failures, context of use, intentions of users or other human actors) (Rismani et al.
+2023a; Weidinger et al.
+2023).
+In this way, safety considerations are reciprocally shaped by social ideas and dynamics that are encoded into AI systems (Dobbe 2022; Rismani et al.
+2023b).
+This concept has been described by Ruha Benjamin (2019) as the “duplicity of technology,” in which AI systems often default to or replicate social power dynamics.
 Addressing this duplicity requires practitioners to understand social inequalities and how they may manifest in sociotechnical systems.
-Without intentional consideration and intervention, T2I systems can easily amplify harmful representations Qadri et al.
-(2023); Bianchi et al.
-(2023).
-2.1.1 AI Governance and Mitigating “Unsafe" Content
-An effective governance approach to cultivating “safe" AI systems is to develop clear safety requirements Leveson (2020), with attention to the sociotechnical nature of AI systems Dobbe (2022) and potential harms Shelby et al.
-(2023).
-There is increasing recognition of representational harms in T2I systems Cho et al.
-(2023); Díaz et al.
-(2023); Bird et al.
-(2023), which are the socially constructed beliefs about different social groups that reinforce unjust hierarchies Wang et al.
-(2022); Barocas et al.
-(2017, 2019), and manifest in different ways (see: Katzman et al.
-(2021)).
-Representational harms are sociotechnical in nature Shelby et al.
-(2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations Birhane et al.
-(2021).
-As such, an increasingly common safety requirement (and social and regulatory expectation) Solaiman (2023); Dias Oliva (2020) is reducing the likelihood that T2I systems will amplify harmful representations — such as demeaning gender, racial, disability, or cultural stereotypes Qadri et al.
-(2023); Bianchi et al.
-(2023); Ungless et al.
-(2023); Naik and Nushi (2023); Bennett et al.
-(2021), particularly when model inputs are benign, unspecified, or do not contain qualifiers (i.e., "irrespective of gender") (Bansal et al., 2022, p.
+Without intentional consideration and intervention, T2I systems can easily amplify harmful representations (Qadri et al.
+2023; Bianchi et al.
+2023).
+AI Governance and Mitigating “Unsafe” Content
+An effective governance approach to cultivating “safe” AI systems is to develop clear safety requirements (Leveson 2020), with attention to the sociotechnical nature of AI systems (Dobbe 2022) and potential harms (Shelby et al.
+2023).
+There is increasing recognition of representational harms in T2I systems (Cho, Zala, and Bansal 2023; Díaz et al.
+2023; Bird, Ungless, and Kasirzadeh 2023), which are the socially constructed beliefs about different social groups that reinforce unjust hierarchies (Wang et al.
+2022; Barocas et al.
+2017; Barocas, Hardt, and Narayanan 2019), and manifest in different ways (Katzman et al.
+2021).
+Representational harms are sociotechnical in nature (Shelby et al.
+2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations (Birhane, Prabhu, and Kahembwe 2021).
+As such, an increasingly common safety requirement (and social and regulatory expectation) (Solaiman 2023; Dias Oliva 2020) is reducing the likelihood that T2I systems will amplify harmful representations — such as demeaning gender, racial, disability, or cultural stereotypes (Qadri et al.
+2023; Bianchi et al.
+2023; Ungless, Ross, and Lauscher 2023; Naik and Nushi 2023; Bennett et al.
+2021), particularly when model inputs are benign, unspecified, or do not contain qualifiers (i.e., “irrespective of gender”) (Bansal et al.
+2022, p.
 1359).
-It is necessary to address representational harms as they perpetuate normative narratives that adversely shape people’s sense of identity and belonging (e.g., Bennett and Keyes (2020); Costanza-Chock (2018); Mengesha et al.
-(2021)) and algorithmic systems can scale the visibility of harmful representations reinforcing their social power Phillips (2018).
-2.1.2 Gender, Social Inequality, and Representation
-Gender inequalities are a persistent global problem, which can be exacerbated by AI technologies Lütz (2023).
-Gender is a social construct through which a society defines expectations around masculinity and femininity, or outside this binary Lorber (1996); West and Zimmerman (1987); West and Fenstermaker (1995).
-As a social construction, gender ideologies are created and changed to meet social, political, and economic needs Martin (2004).
-They are an enduring major component structuring individual, interactional, and institutional power relations Risman (2004).
-A dominant gender construct is the sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment Fredrickson and Roberts (1997); Szymanski et al.
-(2011).
-Scholars have long highlighted how constructed ideas about gender difference are used to justify stratification and inequality Lorber (1994), and that gender ideologies intersect with other axes of inequality, including race and ethnicity, socioeconomic status, sexuality, and disability among others.
-The concept of intersectionality was developed by feminist scholars of color to contest essentialized, white, middle-class, Western constructions of gender (e.g., Anzaldúa (1987); Crenshaw (1991b, a); King (1988); Mohanty (2003)), revealing how gendered representations are informed by intersecting gender, race, and social class ideologies within a “matrix of domination" Collins (1990).
-Computing scholars have examined gendered stereotypes, tropes, and patterns of over/under representation in algorithms Buolamwini and Gebru (2018), datasets Zhao et al.
-(2017); Meister et al.
-(2023); Garrido-Muñoz et al.
-(2023), and interfaces Kay et al.
-(2015); Krieg et al.
-(2023), demonstrating persistent stereotype exaggeration Kay et al.
-(2015).
-Gendered stereotypes affect the performance of real world applications, including online advertising Sweeney (2013), toxic language detection Park et al.
-(2018), machine translation and language technologies Font and Costa-Jussà (2019); Dev et al.
-(2021); Vanmassenhove et al.
-(2018), search engines Kopeinik et al.
-(2023); Kay et al.
-(2015); Albawardi and Jones (2023), and T2I systems Bianchi et al.
-(2023); Naik and Nushi (2023).
-2.2 Stereotypes & Amplification of Societal Biases in T2I Systems
-The representational harm phenomenon in T2I models has been clearly identified (e.g., Hao et al.
-(2023); Bird et al.
-(2023); Bianchi et al.
-(2023); Cho et al.
-(2023); Ungless et al.
-(2023); Wolfe et al.
-(2023); Yu et al.
-(2022); Jain et al.
-(2021); Choi et al.
-(2020)).
-Existing work often focuses on two areas: (1) identifying the kinds and types of harmful representations across different social attributes and axes of identity, and (2) evaluating and mitigating representational societal harms through measurement and technical intervention.
-2.2.1 Amplifying Harmful Societal Representations
-Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone Jain et al.
-(2021); Choi et al.
-(2020).
+It is necessary to address representational harms as they perpetuate normative narratives that adversely shape people’s sense of identity and belonging (e.g., (Bennett and Keyes 2020; Costanza-Chock 2018; Mengesha et al.
+2021)) and algorithmic systems can scale the visibility of harmful representations reinforcing their social power (Phillips 2018).
+Gender, Social Inequality, and Representation
+Gender inequalities are a persistent global problem, which can be exacerbated by AI technologies (Lütz 2023).
+Gender is a social construct through which a society defines expectations around masculinity and femininity, or outside this binary (Lorber 1996; West and Zimmerman 1987; West and Fenstermaker 1995).
+As a social construction, gender ideologies are created and changed to meet social, political, and economic needs (Martin 2004).
+They are an enduring major component structuring individual, interactional, and institutional power relations (Risman 2004).
+A dominant gender construct is the sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011).
+Scholars have long highlighted how constructed ideas about gender difference are used to justify stratification and inequality (Lorber 1994), and that gender ideologies intersect with other axes of inequality, including race and ethnicity, socioeconomic status, sexuality, and disability among others.
+The concept of intersectionality was developed by feminist scholars of color to contest essentialized, white, middle-class, Western constructions of gender (e.g., (Anzaldúa 1987; Crenshaw 1991a, b; King 1988; Mohanty 2003)), revealing how gendered representations are informed by intersecting gender, race, and social class ideologies within a “matrix of domination” (Collins 1990).
+Computing scholars have examined gendered stereotypes, tropes, and patterns of over/under representation in algorithms (Buolamwini and Gebru 2018), datasets (Zhao et al.
+2017; Meister et al.
+2023; Garrido-Muñoz, Martínez-Santiago, and Montejo-Ráez 2023), and interfaces (Kay, Matuszek, and Munson 2015; Krieg et al.
+2023), demonstrating persistent stereotype exaggeration (Kay, Matuszek, and Munson 2015).
+Gendered stereotypes affect the performance of real world applications, including online advertising (Sweeney 2013), toxic language detection (Park, Shin, and Fung 2018), machine translation and language technologies (Font and Costa-Jussà 2019; Dev et al.
+2021; Vanmassenhove, Hardmeier, and Way 2018), search engines (Kopeinik et al.
+2023; Kay, Matuszek, and Munson 2015; Albawardi and Jones 2023), and T2I systems (Bianchi et al.
+2023; Naik and Nushi 2023).
+Stereotypes & Amplification of Societal Biases in T2I Systems
+The representational harm phenomenon in T2I models has been clearly identified (Hao et al.
+2023; Bird, Ungless, and Kasirzadeh 2023; Bianchi et al.
+2023; Cho, Zala, and Bansal 2023; Ungless, Ross, and Lauscher 2023; Wolfe et al.
+2023; Yu et al.
+2022; Jain et al.
+2021; Choi et al.
+2020).
+Existing work often focuses on two areas: (1) identifying the kinds and types of harmful representations or stereotypes across different social attributes and axes of identity, and (2) evaluating and mitigating representational societal harms through measurement and technical intervention.
+It is of note that stereotype amplification and harm amplification, although related, differ in key ways.
+Stereotype amplification in AI systems involves the reinforcement of stereotypes related to a group of people and does not necessarily pertain to safety.
+Conversely, harm amplification as we define it, occurs when AI systems exacerbates harms in the context of safety without necessarily targeting specific demographic groups.
+While stereotype amplification has been well studied, the formalization of harm amplification, to our knowledge, is novel.
+Furthermore, our understanding of harm amplification is related, but distinct, from extant work on “bias amplification” examining how patterns of bias appear in training data, and are amplified statistically in model outputs (e.g., (Zhao et al.
+2017; Wang and Russakovsky 2021; Zhao, Andrews, and Xiang 2023; Lloyd 2018; Taori and Hashimoto 2023; Mehrabi et al.
+2021)).
+In this context, “bias amplification” reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics” (Hall et al.
+2022a, p.
+1), often focusing on binary classification models (e.g., (Leino et al.
+2019; Wang et al.
+2022; Hall et al.
+2022b)) and correlations (Zhao et al.
+2017).
+While bias amplification deals with statistical disparities in model outputs across groups, harm amplification encompasses a broader range of negative consequences not necessarily linked to a specific group.
+Reinforcing Harmful Societal Representations
+Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone (Jain et al.
+2021; Choi et al.
+2020).
 Jain et al.
-Jain et al.
-(2021) demonstrate how generated images have a “distributional shift" away from real world data, in which generative AI systems exacerbate extant social inequalities and patterns of over/under-representation.
+(2021) demonstrate how generated images have a “distributional shift” away from real world data, in which generative AI systems exacerbate extant social inequalities and patterns of over/under-representation.
 The study of representational harms in T2I systems has extended beyond the relationships between training data and model outputs, to focus on the relationships between text prompts and resulting imagery.
-Bianchi and colleagues Bianchi et al.
-(2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts.
+Bianchi and colleagues (2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts.
 They show using Stable Diffusion, for example, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p.
@@
 3).
-Moving beyond representations of people, the authors also show that generating images of everyday objects (e.g., “a house") reflects Western and middle-class stereotypes (p.
+Moving beyond representations of people, the authors also show that generating images of everyday objects (e.g., “a house”) reflects Western and middle-class stereotypes (p.
 3).
@@
 Qadri et al.
-(Qadri et al., 2023, p.
-506) examine the cultural limitations of T2I systems by soliciting South Asian community experts to generate text prompts and annotate the resulting imagery, revealing T2I models “can participate in and scale such existing harmful media regimes of representation and amplify experiences of socio-cultural marginalization." Similarly, Naik et al.
-Naik and Nushi (2023) explore exacerbation of social inequalities along axes of (perceived) gender, age, race, and geographical location, showing state-of-the-art T2I systems amplify real world inequities when compared to U.S.
+(2023, p.
+506) examine the cultural limitations of T2I systems by soliciting South Asian community experts to generate text prompts and annotate the resulting imagery, revealing T2I models “can participate in and scale such existing harmful media regimes of representation and amplify experiences of socio-cultural marginalization.” Similarly, Naik et al.
+(2023) explore exacerbation of social inequalities along axes of (perceived) gender, age, race, and geographical location, showing state-of-the-art T2I systems amplify real world inequities when compared to U.S.
 Bureau of Labor Statistics.
-These studies find that explicitly adding identity terms to unspecified prompts (e.g., “female security guard" Naik and Nushi (2023)) does not effectively mitigate the presence of stereotypes in generated imagery Bianchi et al.
-(2023).
-2.2.2 Evaluating Representational Societal Biases
+These studies find that explicitly adding identity terms to unspecified prompts (e.g., “female security guard” (Naik and Nushi 2023)) does not effectively mitigate the presence of stereotypes in generated imagery (Bianchi et al.
+2023).
+Figure 1: Examples of harm amplifictation for images generated with Stable Diffusion 2.1 with the input prompt shown underneath each image for sexually explicit content (A) and violent content (B).
+A gaussian blur or black box was applied on some images to limit exposure of harms to readers.
+Evaluating Representational Societal Biases
 A parallel body of research focuses not just on identifying the presence and kinds of harmful stereotypes that T2I systems can generate, but also evaluating them.
-Luccioni and colleagues Luccioni et al.
-(2023) examine whether representational societal harms in each modality amplify or compound one another by evaluating images generated from two synthetic text prompt datasets: one with gender and ethnicity terms and one with profession identifiers.
+Luccioni and colleagues (2023) examine whether representational societal harms in each modality amplify or compound one another by evaluating images generated from two synthetic text prompt datasets: one with gender and ethnicity terms and one with profession identifiers.
 The authors analyze representations in T2I imagery, assessing the correlations between social attributes in text markers and the image space, and comparing these to U.S.
 labor demographics, finding consistent under-representation of marginalized identities.
-Wang et al.
 Wang et al.
@@
 Perera et al.
-Perera and Patel (2023) explore social bias with respect to gender, race, and age in the diffusion-based face generation models, finding that diffusion models exacerbate patterns of over- and under-representation, while Saravanan et al.
-(2023) conducts an intersectional analysis of racialized gender stereotypes and professions to support evaluation of T2I models.
+(2023) explore social bias with respect to gender, race, and age in the diffusion-based face generation models, finding that diffusion models exacerbate patterns of over- and under-representation, while Saravanan and colleagues (2023) conducts an intersectional analysis of racialized gender stereotypes and professions to support evaluation of T2I models.
 While a growing body of research is exploring methods to measure and evaluate social stereotypes, safety harm amplification has yet to be formally defined.
 Furthermore, a comprehensive examination of harm representation across social groups will further support sociotechnical safety work and contribute to the social and ethical evaluation of T2I systems.
-3 Defining Harm Amplification
-Our paper focuses on examining representational harms, referring to the socially constructed beliefs about different social groups that reinforce unjust hierarchies Wang et al.
-(2022); Barocas et al.
-(2017); Katzman et al.
-(2021); Barocas et al.
-(2019), in a specific set of generative AI: T2I systems.
-First, we define “unsafe" content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements.
-What could be considered “unsafe" or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations Rismani et al.
-(2023a); Dobbe (2022); Leveson (2020).
-In practice, what constitutes “unsafe" or harmful content may be informed by social, regulatory, and use case expectations; and should be clearly defined with respect to the AI system as a whole.
+Defining Harm Amplification
+Our paper focuses on examining unsafe amplifications of harms in a specific set of generative AI: T2I systems.
+First, we define “unsafe” content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements.
+What could be considered “unsafe” or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations (Rismani et al.
+2023a; Dobbe 2022; Leveson 2020).
+In practice, what constitutes “unsafe” or harmful content may be informed by social, regulatory, and use case expectations; and should be clearly defined with respect to the AI system as a whole.
 For illustrative purposes in this paper, we explore two specific kinds of harmful representations to illustrate how harm amplification can be quantified in T2I systems, following the definitions of Hao et al.
-Hao et al.
 (2023):
@@
   • Violent content: Representations of realistic acts of violence, including but not limited to blood, dismemberment of body parts, and/or displays of aggression/force.
-After harms and safety requirements have been established for a specific T2I system and context, harm amplification can be defined as occurring when the output generated image is notably more harmful than the input text prompt.
+After harms and safety requirements have been established for a specific T2I system and context, harm amplification can be defined as occurring when the output generated image is more harmful than the input text prompt.
 In this sense, amplification reflects when a T2I system increases the severity of the defined harms without being explicitly prompted to do so (e.g., sexually objectifying women in generated imagery).
 Thus, the notion of harm amplification we propose is concerned with understanding the relationships between the text input and resulting imagery.
-We show several examples of harm amplification (generated with Stable Diffusion 2.1 Rombach et al.
-(2022)) in Figure 1 for sexually explicit (Fig 1A) and violent content (Fig 1B) identified by our measurement techniques.
-Figure 1: Blurred harm amplification examples for images generated with Stable Diffusion 2.1 with the input prompt shown underneath each image for sexually explicit content (A) and violent content (B).
-A Gaussian blur was applied on all images to limit exposure of harms to readers.
-Our understanding of harm amplification is related, but distinct, from extant work on “bias amplification" examining how patterns of bias appear in training data, and are amplified statistically in model outputs (e.g., Zhao et al.
-(2017); Wang and Russakovsky (2021); Zhao et al.
-(2023); Lloyd (2018); Taori and Hashimoto (2023); Mehrabi et al.
-(2021)).
-In this literature, “bias amplification" reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics" (Hall et al., 2022a, p.
-1), often focusing on binary classification models (e.g., Leino et al.
-(2019); Wang et al.
-(2022); Hall et al.
-(2022b)) and correlations Zhao et al.
-(2017).
-While bias amplification deals with statistical disparities in model outputs across groups, harm amplification encompasses a broader range of negative consequences not necessarily linked to a specific group.
-Our definition details a more general phenomenon observed in generative AI where the harm for the model output may not be inherently harmful in isolation, but can be harmful when taken into the context of the user input.
-4 Methods to Measure Harm Amplification
-Quantifying harm amplification is a crucial component in assessing “safety" in T2I systems with respect to defined unsafe content.
+In this paper, we choose not to articulate a specific threshold for what constitutes as harm amplification as we believe this should be decided within a specific product or technology context with input from a wider range of stakeholders.
+Instead, we offer a methodological framework that demonstrates how to derive thresholds in various deployment scenarios and show in our experimental section how to employ this framework on an example dataset.
+Framework to Identify Harm Amplification
+Quantifying harm amplification is a crucial component in assessing “safety” in T2I systems with respect to defined unsafe content.
 In this section, we introduce three statistical methods designed to measure harm amplification.
 Leveraging safety classifiers and embedding models, our approach aims to understand how the harms of the generated output images are amplified in the context of the user input prompts.
-Each of our methods has strengths and limitations that make them well-suited for specific cases and available resources.
+Each of our methods has strengths and limitations that makes them well-suited for specific cases and available resources.
 This tailored approach ensures that our methods can be used across diverse contexts contributing to a more robust assessment of harm amplification.
-4.1 Method 1: Distribution-Based Thresholds for Discretized Text Harm Scores
+Method 1 is most appropriate when researchers have access to a large amount of generated data and safety classifiers.
+In cases where data availability is limited, but there is access to safety classifiers, Method 2 may be preferable.
+Method 3 is best suited to low resource situations when researchers can only utilize open source models such as CLIP.
+Method 1: Distribution-Based Thresholds for Discretized Text Harm Scores
 The Distribution-Based Thresholds method utilizes (1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and (2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image).
 Consider a set of text harm scores ($T$) and image harm scores ($I$).
-We first discretize $T$ into $n$ even buckets ($B_{0},B_{1},\text{…},B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
+We first discretize $T$ into $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
 $B_{j}=\{t\in T:l_{j}<t\leq l_{j+1}\},\quad j=0,1,...,n$ (1)
@@
 In our experiments, we fit a first-degree polynomial with the corresponding equation to determine the image harm amplification threshold $HarmAmpThresh_{j}$ for each text harm bucket $B_{j}$:
-$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0}$ (2)
-where $\text{where }y_{j}=\mu_{j}+2\cdot\sigma_{j}$
+$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0},$ (2)
+$where\quad y_{j}=\mu_{j}+2\cdot\sigma_{j}$
 For example, consider a T2I prompt-image dataset where we have image harm scores $I$ and text harm scores $T$.
@@
 We then compare the corresponding image harm score with the harm amplification threshold to determine whether harm is amplified (i.e., $i>HarmAmpThresh_{1}$).
-This method of Distribution-Based Thresholds is best suited to situations where text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier).
+As mentioned previously, this method works best when there is access to large amounts of data to calculate statistical measures from.
+Additionally, this method performs well when text and image safety classifier are not well calibrated (i.e., the output harm severity for an image harm classifier does not equate to the severity of harm outputted by the text harm classifier).
 Rather than comparing text and image harm scores directly which requires well-aligned classifiers, we determine whether an image amplifies harm by comparing its harm score to the relevant distribution within its corresponding text harm bucket, avoiding direct text-image harm score comparisons.
-4.2 Method 2: Bucket Flip for Discretized Text and Image Scores
-Utilizing text and image harm safety scores similar to Method 1 (Section 4.1), the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text.
+Method 2: Bucket Flip for Discretized Text and Image Scores
+Utilizing text and image harm safety scores similar to Method 1, the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text.
 Consider a set of text harm scores ($T$) and image harm scores ($I$), assuming text and image harm scores have a similar scale.
 In cases where the scales differ, techniques such as z-score transformations can be applied to align the scales.
-We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\text{…},B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
+We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
 For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$.
-If the image harm bucket $B_{i}$ is higher than the text harm bucket $B_{t}$, then we determine that harm amplification occurred.
+If the image harm bucket is higher than the text harm bucket ($B_{i}>B_{t}$), then we determine that harm amplification occurred.
 For example, suppose we have discretized harm scores into 10 buckets of ($0-0.1,0.1-0.2,\text{…},0.9-1.0$).
@@
 As the image score bucket is higher than text score bucket, we determine that there was a bucket flip and harm amplification has occurred.
-The Bucket Flip method is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores.
+Unlike Method 1, this method does not require any data and is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores.
 In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical approach for identifying harm amplification.
-4.3 Method 3: Image-Text Co-embedding Based Harm Scores
-While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many people may not have the resources to label such data and thus may not be able to train the necessary safety classifiers.
+Method 3: Image-Text Co-embedding Based Harm Scores
+While training dedicated image or text harm classifiers enables the most direct harm amplification measurement methods, many researchers and practitioners may not have the resources to label such data and thus may not be able to train the necessary safety classifiers.
 In this case, existing large pre-trained image-text models can serve as a substitute.
-Large pre-trained image-text models, such as CLIP Radford et al.
-(2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space.
+Large pre-trained image-text models, such as CLIP (Radford et al.
+2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space.
 Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s embedding, $i$, is to a harmful concept in the embedding space than the prompt’s embedding, $p$, is.
-Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the harm sexually explicitly, we can use the words “sexualized", “porn", etc.), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$.
+Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the sexually explicit harm category, we can use the words “sexual”, “porn”, etc.
+- see Appendix, Table 2), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$.
 These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space.
-Harm amplification can then be quantified as $mean(cosine\_similarity(h_{k},i)-cosine\_similarity(h_{k},p))for\,all\,k$.
+We can then calculate the average distance between the proximity of image embeddings and harm concepts compared to prompt embeddings and harm concepts:
+$D_{harm}=\mu(cos\_similarity(h_{k},i)\\ -cos\_similarity(h_{k},p))\mid k\in\{0,1,2,...,K\}$ (3)
+Harm amplification is determined to occur if $D_{harm}$ is greater than some predefined harm amplification threshold.
 This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt.
-5 Experiments
-In this section, we demonstrate how to employ the methods that we outlined in Section 4 on a large and varied T2I measurement dataset designed to simulate usage in real-world model deployment to develop a harm amplification criteria for each method.
-We then evaluate how well our methods perform by applying these criterion on an independent human annotated T2I evaluation dataset by calculating precision, recall, and f1-scores.
-5.1 Measurement Dataset & Models
-Our measurement dataset consists of 497,157 prompts representing various demographics and prompt categories.
+Experiments
+In this section, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset.
+These methods are then evaluated on an independent human annotated dataset where f1-scores, precision, and recall are derived.
+Measurement Dataset and Models
+To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories.
 These prompts were sampled from dogfood user data aimed at testing generative AI models.
@@
 While curated adversarial prompts can be valuable for specific safety assessments, our method takes a broader perspective by aiming to understand how T2I models can generate harm amplifying images over a large range of prompts (from benign to borderline to harmful).
-We used Stable Diffusion 2.1 Rombach et al.
-(2022) to generate images for our prompt measurement dataset.
+We used Stable Diffusion 2.1 (Rombach et al.
+2022) to generate images for our prompt measurement dataset.
 Four images were generated per prompt resulting in 1,988,628 total generated images.
@@
 Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
-Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content.
+Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2.
 Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1.
-Fairness machine learning classifiers that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1.
-For the pre-trained image-text model, we use CLIP Radford et al.
-(2021) and generated embeddings for images and their corresponding prompts.
-This large, varied measurement dataset served as the foundation for applying each of our methods to establish criteria for determining harm amplification in sexually explicit and violent content.
-Machine annotations for sexually explicit content, violent content, as well as CLIP embeddings served as inputs for our measurement techniques.
-After applying each of our measurement techniques to our measurement dataset, a criteria was obtained for each method:
-  1. 1.
-    Distribution-Based Threshold: Harm amplification occurs if the image harm is higher than a threshold for a given text harm bucket.
-  2. 2.
-    Bucket Flip: Harm amplification occurs if the image harm bucket is higher than the corresponding text harm bucket.
-  3. 3.
-    Image-Text Co-embedding: Harm amplification occurs if an image CLIP embedding is relatively closer to the harm concept embedding compared to the prompt embedding.
-5.2 Evaluation Data & Metrics
-To create the evaluation dataset, we sourced 742 prompt, image pairs from the Adversarial Nibbler data challenge Alicia Parish (2023).
-This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 AI (2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) AI (2022).
-At least five raters annotated text and image safety independently as well as identified the type of harm (i.e., sexually explicit or violent) that was in the image if the image was deemed unsafe.
-Majority vote was used to determine the final label for each prompt harm and image harm label.
-We defined ground truth (if harm amplification occurred) as when the text prompt was considered safe by human annotators, but the image was deemed harmful.
-Ground truth for the absence of harm amplification was defined as when human annotators judged the text and image both to be safe.
-In addition to human labels, we obtained CLIP embeddings and machine annotations for sexually explicit and violent content for both prompts and images that served as inputs for each of our methods to infer whether harm amplification occurred.
-Subsequently, using the criteria established for each method (derived from our measurement data), we assessed the presence of harm amplification in the Nibbler evaluation for each harm, using human labels as ground truth.
-We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset.
-This evaluation was conducted separately for sexually explicit and violence harm amplification.
-Furthermore, perceived gender was additionally machine annotated in the evaluation dataset to investigate potential disparities in rates of harm amplification across genders.
-5.3 Experimental Results
-5.3.1 Sexually Explicit & Violent Content Results
-To quantify harm amplification using the Distribution-Based Threshold method (Section 4.1), we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence.
-Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix A, Figure 2A).
-The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket.
-A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix A, Figure 2B) for each harm type.
-For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution.
-Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset.
-Notably, we observed high precision (0.925) and recall (0.973) when evaluating sexually explicit content in the Nibbler dataset (see Table 1).
-However, we noted slightly lower precision (0.788) and recall (0.512) when evaluating this measurement technique on violent content (Table 1).
-Using the Bucket Flip method (Section 4.2), we first calculated and recorded the mean and standard deviation of both the text harm scores and image harm scores for the measurement data for each harm type: sexually explicit and violence.
-We then did z-score transformations on both the text and image harm scores to transform them into similar scale.
-We then subsequently divided that range into 10 even buckets.
-For evaluation on the Nibbler dataset, we similarly applied z-score transformations to the machine annotations for each harm type using the corresponding mean and standard deviation on the text and image derived from the measurement data.
-We then bucketized the transformed scores with the bucket boundaries obtained from the measurement data and assessed whether the image buckets were higher than the text buckets.
-We observed high precision (0.896) and recall (0.956) when evaluating sexually explicit content and slightly lower precision (0.870) and recall (0.488) when evaluating on violent content.
-While we expect the Image-Text Co-embedding method (Section 4.3) to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources.
-We used CLIP to encode images and prompts, and defined the sexually explicit harm concept using the words “porn," “sexual," and “nudity." We defined the violence harm concept using the words “violent," “weapons," and “injury".
-Utilizing more terms to define each harm concept did not have a significant effect on results (see Appendix B, Figure 5).
-Thus, we report results using the specified three terms for each harm concept that we mentioned above.
-Embedding and distance scores were calculated as described in Section 4.3 separately for the measurement data and evaluation data.
-Various harm amplification thresholds were applied (using distances calculated from the measurement data embeddings) to the evaluation dataset to obtain precision-recall curves (see Appendix B, Figure 6).
-Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case.
-For these results we set an operating recall of 0.6, and compute other precision and f1-scores from this number using the PR curves.
-As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.821, recall of 0.600) than violence harm amplification (precision 0.390, recall 0.600).
+For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al.
+2021).
+In addition to safety classfiers and image-text embedding models, fairness machine learning classifiers (Baruah et al.
+2022) that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1 in the evaluation data.
 Harm Amplification Measurement Method Sexually Explicit Violence
 Precision Recall F1-Score Precision Recall F1-Score
-Distribution-Based Thresholds 0.925 0.973 0.948 0.788 0.512 0.621
-Bucket Flip 0.896 0.956 0.925 0.870 0.488 0.625
-Image-Text Co-embedding 0.821 0.600 0.693 0.390 0.600 0.473
+Distribution-Based Thresholds 0.864 0.985 0.920 0.460 0.831 0.592
+Bucket Flip 0.910 0.950 0.930 0.708 0.554 0.622
+Image-Text Co-embedding 0.585 0.920 0.715 0.322 0.566 0.410
 Table 1: Harm Amplification measurement methods evaluation on Nibbler data.
@@
 Precision, recall, and F1-scores for sexually explicit content (left 3 columns) and violence (right 3 columns) were reported for each of the three measurement methods.
-5.3.2 Gender Fairness
-As described in Related Work (Sections 2.1.2 & 2.2.1), the amplification of social stereotypes disproportionately affects marginalized communities Bianchi et al.
-(2023); Naik and Nushi (2023).
-We present quantitative measurement of harm amplification — particularly the gendered sexual objectification stereotype Fredrickson and Roberts (1997); Szymanski et al.
-(2011) — and show that it disproportionately affects images depicting perceived females.
-We recognize the dynamic and non-binary nature of gender; however, in this preliminary study, we focus on perceived binary (male/female) gender expressions due to the ease in methodology.
-We also acknowledge that we that cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
-To show the prevalence of disparity in harm amplification across perceived gender expression in Nibbler dataset for the harm types of sexual explicit and violent content, we evaluate the efficacy of our proposed methods across the gender demographic on the Nibbler dataset.
-From the Nibbler dataset, for all images, both the number of people and their perceived gender expression were determined by machine classification Baruah et al.
-(2022).
-We annotate the tuples in which the number of perceived females are in strict majority to be a (perceived) female data point, and do the same for (perceived) male gender.
-We do not include any tuple in which, either there are no faces or in which the number of perceived females are equal to the number of perceived males.
-Gender Expression Sexually Explicit Violence
-# Text Safe Examples Harm Amplification Rate # Text Safe Examples Harm Amplification Rate
-(Perceived) Female 173 0.80 60 0.43
-(Perceived) Male 145 0.46 120 0.36
-Table 2: Harm Amplification rates across perceived gender expressions on Nibbler data. Significance testing was done using z-test. Statistically significant numbers are in bold ($p<0.01$).
-We calculate the rates of harm amplification across both the perceived gender expressions using the ground truth of the Nibbler dataset for both harm categories described in Section 3.
-Table 2 shows the the number of examples belonging to each category across the two perceived gender expressions.
-We find the harm amplification rate for (perceived) females is statistically significantly higher ($p<0.01$) for the category of sexually explicit harm amplification.
-This important observation quantifies the increased sexually explicit harm amplification present in T2I systems for the perceived female gender.
-Furthermore, the increased rate of harm amplification of violent content is not statistically significant ($p=0.181$).
-A good method for detecting harm amplification should run equally well across all the demographic groups.
-The Distribution-Based Thresholds provide comparable efficacy in detection of harm amplification instances for sexually explicit content across both the perceived gender expressions (see Appendix C, Table 3).
-For the remaining methods, the precision in detecting over-sexualization in perceived females is higher than the precision in detecting it in perceived males.
-6 Discussion & Conclusion
-In this paper, we provide a formal definition of harm amplification in T2I models as occurring when the generated output image reflects more harmful representations that were not explicit in the text input (Section 3).
-Using two specific kinds of harmful representations (sexually explicit content and violent content), we showed how to quantify harm amplification by providing three methods that use safety classifier harm scores and CLIP embeddings (Section 4).
-We then empirically illustrated the application of these methodologies on a measurements dataset that simulated real-world model deployment and evaluated these methods on an independent human annotated dataset (Section 5).
-6.1 Using Data-Driven Methods to Quantify Harm Amplification
-Each of our three data-driven methods for quantifying harm amplification offers its own strengths and limitations.
-The Distribution-Based Thresholds method may be most appropriate if text and image harm scores are not well calibrated.
-However, with this method, we assume that image harm scores within each bucket follow a normal distribution.
-If this assumption is violated, then thresholds for each bucket may not be effective in quantifying when harm amplification occurs.
-The Bucket Flip method is the easiest method to employ, but best performance is achieved if the text and image harm scores are well calibrated – otherwise, a bucket flip may not necessarily mean that harm is amplified.
-Another limitation for this method is that (text, image) pairs with close harm scores may be positioned in different buckets which might be assessed as harm amplification instances even when the actual difference in scores is small.
-While the Image-Text Co-embedding method performs well in a low-resource setting, a fine-tuned model using CLIP embeddings specifically trained to detect harm is likely to yield better results.
-As we observe in the methods using safety classifiers, specific use-case models can dramatically improve performance when detecting instances of harm amplification.
-Across all three methods, we assume the harm amplification criterion derived from the measurements dataset will generalize to new data generated by our measurement model (Stable Diffusion 2.1) as well as new data generated by other T2I models.
-This assumption holds if there is similar harm amplification rates across different T2I models.
-Although we observe good performance in our evaluation dataset that encompasses data generated from several different models, further research should explore the generalizability of harm amplification methods in cross-domain T2I models.
-6.2 Reduced Reliance on Human Annotation
-Our novel framework significantly reduces the reliance on human annotations, which may be especially helpful in low-resource settings.
+Evaluation Data & Metrics
+To create the evaluation dataset, we sourced 1125 prompt, image pairs from the Adversarial Nibbler data challenge (Quaye et al.
+2024).
+This public, crowd-sourced challenge aimed to identify failure-modes of DALL-E 2 (Open AI 2023) and various Stable Diffusion versions (XL 1.0, 1.5, and 2.1) (Stability AI 2022).
+Five to six raters trained on AI safety policies annotated prompts and images separately for the presence of sexually explicit content or violent content.
+To derive a confidence score for sexually explicit and violence, we converted the ratings of the annotators into a proportional measure (i.e., if 3 out of 5 raters rated an image as violent, the violence score for that image would be 0.6 or 60%).
+The ground truth labels for harm amplification for each harm were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm).
+If the image confidence score was less than or equal to the corresponding text confidence score for a given harm, then the ground truth was labeled as harm amplification not occurring.
+We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset.
+This evaluation was conducted separately for sexually explicit and violence harm amplification.
+Furthermore, we conducted an additional analysis on the evaluation dataset to examine gender fairness.
+Specifically, we investigated whether harm amplification of sexually explicit or violent content disproportionately affected one perceived gender.
+Perceived gender was machine annotated in the evaluation dataset with significance testing performed on rates of harm amplification across genders.
+Experimental Results
+Evaluation of Methods in Detecting Harm Amp
+To quantify harm amplification using Method 1, Distribution-Based Threshold Method, we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence.
+Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix, Figure 3A).
+The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket.
+A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 3B) for each harm type.
+For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution.
+Having derived fitted image harm amplification thresholds for each text bucket for each harm type, we then assessed how well our measurement technique performed on the Nibbler evaluation dataset.
+A precision of 0.864 and recall of 0.985 was observed when evaluating sexually explicit content in the Nibbler dataset (see Table 1).
+When evaluating violent harm amplification, we noted a precision of 0.460 and recall of 0.831.
+Using Method 2, the Bucket Flip Method, we divided text and image harm scores obtained by safety classifiers applied on the Nibbler data into 5 even buckets.
+We then assessed for each prompt, image pair whether the image harm bucket was larger than the text harm bucket.
+This approach yielded better results than Method 1 with a precision of 0.910 and recall of 0.950 for sexually explicit harm amplification, and a precision of 0.708 and recall of 0.554 for violence harm amplification.
+While we expect Method 3, Image-Text Co-embedding Method, to perform worse at the task of identifying instances of harm amplification than using dedicated sexually explicit and violence text and image classifiers, this method is useful for those with access to limited resources.
+We used CLIP to encode images and prompts, and defined sexually explicit harm concept using 15 words such as “porn”, “sexual”, and “nude” (see Appendix, Table 2).
+These concept words were sourced from internal adversarial testing and (Rando et al.
+2022a) in their attempt to reverse-engineering Stable Diffusion’s NSFW filter.
+Adversarial testing was used to find 15 violence concept words such as “violence”, “weapons”, and “blood” (see Appendix, Table 2).
+Various harm amplification thresholds were applied to the evaluation dataset to obtain precision-recall curves (see Appendix, Figure 5).
+Unlike the other methods, there is not an obvious way to define what the best harm amplification threshold is, but rather, practitioners can decide on an appropriate threshold based on their specific use case.
+For these results we obtained a threshold that maximized the f1 score, and computed precision and recall from this number using the PR curves.
+As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.585, recall of 0.920) than violence harm amplification (precision 0.322, recall 0.566).
+Gender Disparities in Harm Amplification
+As described in the Related Works section, the amplification of social stereotypes disproportionately affects marginalized communities (Bianchi et al.
+2023; Naik and Nushi 2023).
+In the evaluation dataset, we analyzed ground truth harm amplification rates and their disparate impacts on perceived genders.
+In this study, we recognize the dynamic and non-binary nature of gender, but ultimately chose to focus on perceived binary (male/female) gender expressions due to the ease in methodology.
+We also acknowledge that we cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
+Machine annotations of perceived gender (Baruah et al.
+2022) were collected for each image in the Nibbler evaluation dataset.
+We restricted our analysis to images in which there was a strict majority of a perceived gender.
+Images in which there were no faces or where the number of perceived males equaled that of perceved females were discarded.
+We then calculated the rates of harm amplification using the majority perceived gender.
+Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2), in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011).
+There was no significant difference in violence harm amplification across perceived genders ($p=0.492$).
+We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
+Figure 2: Difference in harm amplification rates across perceived genders in the Nibbler evaluation dataset.
+Images containing females were significantly more oversexualized than males whereas there was no significant difference in harm amplification of violence.
+$*p<0.05,\ **p<0.01,\ ***p<0.001$
+Limitations
+While this framework is shown to be effective in identifying harm amplification, we acknowledge the limitations of our methodologies and identify potential areas of future research.
+Method 1, the Distribution-Based Thresholds Method is highly sensitive to the measurement dataset’s distribution of scores.
+If the distribution is skewed (i.e., the dataset is predominantly either safer or more harmful), derived harm amplification thresholds will then be biased.
+Thus, when using this method, it is imperative to analyze and verify that the distribution of harm scores within the measurement dataset aligns with practitioners’ assumptions.
+As noted previously, Method 2, the Bucket Flip Method relies on the successful calibration of text and image safety classifiers.
+Uncalibrated classifiers can result in inaccurate identification of harm amplification.
+Future research should investigate how calibration methods and various types of safety classifiers affect this method.
+Method 3, the Image-Text Co-embedding Method, utilized the open source model CLIP to find the difference in embeddings of images and prompts in relation to harm concept words.
+However, CLIP has been shown to exhibit fairness issues particularly concerning demographic attributes such as gender and race (Luo et al.
+2024; Wang, Zhang, and Sang 2022).
+A potential future direction of work could involve using a debiased version of CLIP similar to that of FairCLIP (Wang, Zhang, and Sang 2022) to investigate its efficacy in identifying harm amplification instances.
+Across all methods, we see higher performance in identifying instances of harm amplification in sexually explicit content compared to violent content.
+Further research should investigate the factors contributing to this disparity in performance such as differences in safety classifier effectiveness, noise in human annotations, or biases within datasets used.
+Discussion
+In this paper, we provide the first formalization of harm amplification in T2I models, defining it as occurring when the generated output image reflects more harmful representations that were not explicit in the text input.
+Using two specific kinds of harmful representations (sexually explicit content and violent content), we provided a framework of different methods to identify instances of harm amplification under different resource and deployment scenarios.
+We then empirically illustrated the application of this framework on an independent human annotated dataset.
+Reduced Reliance on Human Annotation
+Our framework significantly reduces the reliance on human annotations.
 By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets.
@@
 Thus, our proposed framework not only offers robust methodologies for measuring harm amplification but also addresses practical challenges that researchers may encounter when deploying models at scale.
-Of course, machine annotations — akin to human annotations Davani et al.
-(2022); Denton et al.
-(2021) — can also reflect social biases.
+Of course, machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al.
+2021) — can also reflect social biases.
 These biases may stem from the subjective nature of defining ground truth in safety.
 Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
-6.3 Gender Disparities in Harm Amplification
+Gender and Harm Amplification
 In analyzing rates of harm amplification in the evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content.
@@
 Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts in different sociodemographic groups.
-By identifying gender related disparities, we contribute to existing work Luccioni et al.
-(2023); Naik and Nushi (2023) demonstrating gender stereotypes in T2I models.
-The durability of gender stereotypes in AI systems Buolamwini and Gebru (2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes.
+By identifying gender related disparities, we contribute to existing work (Luccioni et al.
+2023; Naik and Nushi 2023) demonstrating gender stereotypes in T2I models.
+The durability of gender stereotypes in AI systems (Buolamwini and Gebru 2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes.
 While we focused on gender, other axes of inequality, such as race, ethnicity, disability, sexuality, and culture, among others, should be explored.
-6.4 Future Work
+Future Work
 While our work lays a foundation to better understand and quantify harm amplification in T2I systems, further research is needed to more holistically understand the phenomenon and develop strategies for its reduction in T2I models.
@@
 The development of these mitigation techniques will contribute to the overall responsible deployment of T2I systems, supporting the increased protection and safety of users.
-7 Acknowledgements
-We thank Kathy Meier-Hellstern, Alexander Vorontsov, and Vinodkumar Prabhakaran for their feedback and suggestions.
-We also thank Auriel Wright, Candice Schumann, Lora Aroyo, and Alicia Parrish for discussions on analysis and writing.
-Appendix A Experimental Methods Applied on the Measurement Dataset
+Appendix A Appendix
 Figure 3: Method 1: Distribution-based thresholds applied on the measurements dataset.
@@
 B) We then repeated the same procedure for machine annotations for violence.
-Appendix B Evaluation Results on Nibbler Data
-Figure 3: Confusion Matrix for Method 1: Distribution-based thresholds on the Nibbler evaluation dataset.
-Using the thresholds for this method applied on the measurement dataset, we evaluated these thresholds on the Nibbler data with human annotations serving as ground truth.
-In the left figure, we see the confusion matrix for harm amplification predictions for sexually explicit content in Nibbler (negative being no harm amplification occurred and positive being that harm amplification occurred).
-In the right figure, we show the confusion matrix for harm amplification predictions for violent content in Nibbler.
-Figure 4: Confusion Matrix for Method 2: Bucket Flip for for Text and Image Score Bucket on the Nibbler evaluation dataset.
-With this method applied on the measurement dataset, we evaluated the bucket flip criteria on the Nibbler data with human annotations serving as ground truth.
-The left figure shows the confusion matrix for sexually explicit content and the right figure shows the confusion matrix for violent content
-Figure 5: Comparing the PR curves with different amount of terms in Method 3.
-For sexually explicit content, we show no notable difference between the PR curves for 3 terms ("porn", "sexual", "nudity") and 10 terms ("porn", "sexual", "nudity", "erotic", "sex", "breasts", "penis", "vagina", "erotica", and "indecency").
-AUC difference was also minimal (3 terms: 0.846, 10 terms: 0.848).
-For violent content, we again see minimal difference between the PR curves for 3 terms ("violent", "weapons", "injury") and 10 terms ("violent", "weapons", "injury", "gore", "blood", "death", "gruesome", "pain", "wound", "brutality").
-Again, AUC difference between 3 terms (0.412) and 10 terms (0.397) was small.
-Figure 6: Precision-recall curves for the Image-Text Co-embedding method for sexually explicit content (left) and violent content (right) for the final 3 terms used.
-Appendix C Gender Fairness Analysis
-Table 3 shows the efficacy of our proposed methods across both the perceived gender expressions for the harm amplification of sexually explicit content.
-Please note, we focus on sexually explicit harm only as the harm amplification rates were not statistically different for violent content across the two perceived gender expressions.
+Sexually Explicit
+porn sexual
+nude sex
+18+ naked
+nsfw dick
+vagina explicit content
+uncensored fuck
+nipples naked person
+sexy
+Violence
+violence gore
+blood attack
+bloodshed war
+horror fight
+weapons injury
+death pain
+wound brutality
+harm
+Table 2: Harm concepts for Method 3: Image-Text Co-embedding. These harm concepts were meant to approximate the embedding of sexually explicit and violence.
+Figure 4: Confusion matrix for each method (left: Distribution-Based Thresholds, middle: Bucket Flip, right: Image-Text Co-embedding) evaluated on the Nibbler dataset.
+Figure 5: Precision-recall curves for the Image-Text Co-embedding method for sexually explicit content (left) and violent content (right).
 Harm Amplification Measurement Method Female Male
 Precision Recall F1-Score Precision Recall F1-Score
-Distribution-Based Thresholds 0.978 0.971 0.974 0.970 0.970 0.970
-Bucket Flip 0.985 0.964 0.974 0.753 0.865 0.805
-Image-Text Co-embedding 0.978 0.619 0.758 0.760 0.567 0.650
+Distribution-Based Thresholds 0.858 0.979 0.915 0.792 0.987 0.878
+Bucket Flip 0.913 0.932 0.923 0.860 0.961 0.908
+Image-Text Co-embedding 0.656 0.911 0.763 0.308 0.844 0.451
 Table 3: Efficacy of the proposed methods across the two perceived gender expressions for sexually explicit harm amplification.

### v3

@@
 Text-to-image (T2I) models have emerged as a significant advancement in generative AI; however, there exist safety concerns regarding their potential to produce harmful image outputs even when users input seemingly safe prompts.
-This phenomenon, where T2I models generate harmful representations that were not explicit in the input, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms.
+This phenomenon, where T2I models generate harmful representations that were not explicit in the input prompt, poses a potentially greater risk than adversarial prompts, leaving users unintentionally exposed to harms.
 Our paper addresses this issue by formalizing a definition for this phenomenon which we term harm amplification.
@@
 We then empirically examine how to apply these different methodologies to simulate real-world deployment scenarios including a quantification of disparate impacts across genders resulting from harm amplification.
+Additionally, we employ explainability techniques to understand mechanisms driving harm amplification and how specific input elements contribute to harmful outputs.
 Together, our work aims to offer researchers tools to comprehensively address safety challenges in T2I systems and contribute to the responsible deployment of generative AI models.
 Introduction
-Generative text-to-image (T2I) systems allow users to create new image content in response to a text input, learning patterns and relationships from large-scale training data (Open AI 2023; Midjourney Inc.
+Generative text-to-image (T2I) systems allow users to create new image content in response to a text input prompt.
+These systems learn patterns and relationships from their large-scale training data (Open AI 2023; Midjourney Inc.
 2022; Stability AI 2022).
-As all data are reflections of the social world (Gitelman 2013), training data can mirror the harmful stereotypes and social inequalities that exist offline (Birhane et al.
+However, they frequently reflect harmful stereotypes and social inequalities that are embedded in their training data (Birhane et al.
 2023; Birhane, Prabhu, and Kahembwe 2021; Kirk et al.
 2021; Wang et al.
-2022), which are subsequently replicated in generated imagery (Cho, Zala, and Bansal 2023).
-Responsible AI scholarship illustrates how content from T2I systems can reflect and amplify the harmful representations or stereotypes of groups that exist offline (Qu et al.
+2022), which are subsequently reproduced in generated imagery (Cho, Zala, and Bansal 2023).
+The interaction between these technical systems and existing social power dynamics in the world (Shelby et al.
+2023; Green and Viljoen 2020) can further perpetuate and amplify the harmful representations of marginalized groups (Qu et al.
 2023; Qadri et al.
@@
 2022b).
-Harms from generative AI — like all algorithmic systems — occur from the interplay of technical system components and social power dynamics (Shelby et al.
-2023; Green and Viljoen 2020).
-Thus, effective harm reduction work requires practitioners to develop sociotechnical safety requirements, conduct evaluations, and implement strategies to proactively minimize harmful representations prior to launch (Rismani et al.
-2023a; Dobbe 2022; Solaiman et al.
-2023).
-Responsible AI research has cultivated safety techniques to curb the generation of unsafe content including through reinforcement learning (Daniels-Koch and Freedman 2022), dataset remediation (Birhane et al.
-2023), red teaming (Parrish et al.
+Responsible AI research has often approached the study of harmful content in generative AI systems using adversarial prompts and red teaming (Parrish et al.
 2023; Rando et al.
-2022b), model input and output filters (Hao et al.
-2023), and techniques to suppress harmful content generation (Schramowski et al.
-2023).
-These component safety techniques play a vital role in reducing the likelihood a model will generate harmful representations.
-However, these techniques do now eliminate all potential harms such as the exposure of unintentional contextualized harms in generative AI systems.
-Consider a scenario where a user requests an image of “people at a party” from a multimodal AI model, and images of women in bikinis dancing are outputted.
-The input prompt and output image in isolation may be considered safe, but when taken together in context of each other, they can expose unintentional harm (oversexualization of women) to the user.
-This scenario underscores the complex challenges posed by multimodal systems.
-Harmful representations may result from computational, contextual, and compositional risks between text-image pairings with visual content introducing “additional shock factor[s]” (Weidinger et al.
-2023, p.
-13).
+2022b).
+While these methods are valuable, they overlook a critical aspect of safety - unintentional contextual harms arising from seemingly benign inputs.
+For example, a user requesting an image of a “black gay man” may inadvertantly receive sexualized images from a generative AI system, exposing them to unintentional harm through oversexualization which may further perpetuate harmful societal biases.
+Even with safety filters in place (Hao et al.
+2023), users may experience frustration and distress when safe prompts are blocked without explanation.
+This scenario underscores the complex challenges posed by multimodal systems where harmful representations can result from computational, contextual, and compositional risks.
 While responsible AI scholarship acknowledges the amplification of harm as a key negative consequence of unsafe AI systems, including the “prospect of algorithmic systems exacerbating or scaling existing social inequalities” (Shelby et al.
 2023, p.
-733), a precise definition of multimodal safety harm amplification has yet to be introduced, to our knowledge.
-To address these challenges, we formalize the concept of harm amplification in T2I systems and present a framework to quantify harm amplification.
-This framework employs three statistical methods that can be used in different contexts, including scenarios with: 1) large amounts of data and access to safety classifiers 2) limited data, but having access to safety classifiers, and 3) low resource settings with access only to open source models.
-In summary, we make the following research contributions:
+733), there is currently no precise definition of multimodal safety harm amplification.
+To address these challenges, we provide the first formalization of harm amplification and make the following research contributions:
   • A formal definition of harm amplification for T2I models as occurring when the generated image from a T2I model reflects harmful or unsafe representations that were not explicit in the text input.
   • A framework that includes three methods for quantifying and measuring harm amplification instances in T2I models.
-These methods include using other pretrained models such as safety classifiers and CLIP (Radford et al.
-2021) to determine whether harm amplification has occurred (i.e., when the output image is more harmful than the user input prompt).
-  • An empirical examination of how to employ each methodology to simulate different real-world model deployment scenarios, which can dynamically inform practitioners’ safety frameworks.
-This includes a quantification of disparate impact across gendered outputs as a result of harm amplification.
-In what follows, we first orient our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification; followed by a description of our novel methodologies.
-We then detail our experiments and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems, arguing for greater attention to harm amplification in responsible AI research and practice.
+  • An empirical examination of patterns in harm amplification including quantifying its disparate impact across genders and applying explainability techniques to identify aspects of the user input that contribute to harmful output generation.
+In what follows, we first orient our work with respect to key concepts, including sociotechnical approaches to safety in T2I systems and stereotype amplification; followed by a description of our novel methodologies and their efficacy in measuring harm amplification.
+We then examine patterns of harm amplification within our data and finish with a discussion of how empirically measuring harm amplification strengthens safety work on T2I systems, arguing for greater attention to harm amplification in responsible AI research and practice.
 Related Work
-This research lies at the intersection and builds on responsible AI research on harm, safety, and multimodal generative AI.
-The contributions of this research extend the work summarized, and highlight the need for a formalized definition of- and evaluating- harm amplification.
+This research builds on existing responsible AI literature addressing harm, safety, and multimodal generative AI, specifically T2I systems.
+Our contributions extend the work summarized, and highlight the need for a formalized definition of harm amplification and methods for evaluating it.
 Sociotechnical Safety in T2I Systems and Harm Reduction
-Concerns about safety in AI systems encompass a range of diverse sociotechnical issues that directly or indirectly shape the well-being of people or the environment.
-As such, safety is contingent and contextual, and an emergent characteristic or property of AI systems (Leveson 2020).
-While approaches to safety in AI systems vary (Ahmed et al.
-2023), a key concern of a systems approach is to identify system-level hazards (Dobbe 2022) and integrate safety considerations into the development process (Leveson 2016) to minimize the likelihood that harms will arise from system use or failure (Dobbe 2022).
-Within this approach, social and ethical safety considerations of T2I systems need to be identified by proactively examining the interaction between components of the AI system (e.g., technical capabilities, potential failures, context of use, intentions of users or other human actors) (Rismani et al.
-2023a; Weidinger et al.
+AI safety encompass a range of diverse sociotechnical issues that impact the well-being of people and the environment.
+A systems approach to AI safety identifies hazards (Dobbe 2022) and integrates safety considerations into the development process (Leveson 2016) to minimize potential harms (Dobbe 2022).
+Addressing social and ethical safety in T2I systems requires examining the interactions between AI components and social dynamics (Rismani et al.
+2023; Weidinger et al.
 2023).
-In this way, safety considerations are reciprocally shaped by social ideas and dynamics that are encoded into AI systems (Dobbe 2022; Rismani et al.
-2023b).
-This concept has been described by Ruha Benjamin (2019) as the “duplicity of technology,” in which AI systems often default to or replicate social power dynamics.
-Addressing this duplicity requires practitioners to understand social inequalities and how they may manifest in sociotechnical systems.
-Without intentional consideration and intervention, T2I systems can easily amplify harmful representations (Qadri et al.
+AI systems can replicate social dynamics (Benjamin 2019) or even amplify harmful representations that perpetuate societal biases (Qadri et al.
 2023; Bianchi et al.
 2023).
-AI Governance and Mitigating “Unsafe” Content
 An effective governance approach to cultivating “safe” AI systems is to develop clear safety requirements (Leveson 2020), with attention to the sociotechnical nature of AI systems (Dobbe 2022) and potential harms (Shelby et al.
@@
 2023), as social beliefs about people, culture, and experiences are encoded into systems through training data and learned associations (Birhane, Prabhu, and Kahembwe 2021).
-As such, an increasingly common safety requirement (and social and regulatory expectation) (Solaiman 2023; Dias Oliva 2020) is reducing the likelihood that T2I systems will amplify harmful representations — such as demeaning gender, racial, disability, or cultural stereotypes (Qadri et al.
-2023; Bianchi et al.
-2023; Ungless, Ross, and Lauscher 2023; Naik and Nushi 2023; Bennett et al.
-2021), particularly when model inputs are benign, unspecified, or do not contain qualifiers (i.e., “irrespective of gender”) (Bansal et al.
-2022, p.
-1359).
-It is necessary to address representational harms as they perpetuate normative narratives that adversely shape people’s sense of identity and belonging (e.g., (Bennett and Keyes 2020; Costanza-Chock 2018; Mengesha et al.
-2021)) and algorithmic systems can scale the visibility of harmful representations reinforcing their social power (Phillips 2018).
-Gender, Social Inequality, and Representation
-Gender inequalities are a persistent global problem, which can be exacerbated by AI technologies (Lütz 2023).
-Gender is a social construct through which a society defines expectations around masculinity and femininity, or outside this binary (Lorber 1996; West and Zimmerman 1987; West and Fenstermaker 1995).
-As a social construction, gender ideologies are created and changed to meet social, political, and economic needs (Martin 2004).
-They are an enduring major component structuring individual, interactional, and institutional power relations (Risman 2004).
-A dominant gender construct is the sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011).
-Scholars have long highlighted how constructed ideas about gender difference are used to justify stratification and inequality (Lorber 1994), and that gender ideologies intersect with other axes of inequality, including race and ethnicity, socioeconomic status, sexuality, and disability among others.
-The concept of intersectionality was developed by feminist scholars of color to contest essentialized, white, middle-class, Western constructions of gender (e.g., (Anzaldúa 1987; Crenshaw 1991a, b; King 1988; Mohanty 2003)), revealing how gendered representations are informed by intersecting gender, race, and social class ideologies within a “matrix of domination” (Collins 1990).
-Computing scholars have examined gendered stereotypes, tropes, and patterns of over/under representation in algorithms (Buolamwini and Gebru 2018), datasets (Zhao et al.
-2017; Meister et al.
-2023; Garrido-Muñoz, Martínez-Santiago, and Montejo-Ráez 2023), and interfaces (Kay, Matuszek, and Munson 2015; Krieg et al.
-2023), demonstrating persistent stereotype exaggeration (Kay, Matuszek, and Munson 2015).
-Gendered stereotypes affect the performance of real world applications, including online advertising (Sweeney 2013), toxic language detection (Park, Shin, and Fung 2018), machine translation and language technologies (Font and Costa-Jussà 2019; Dev et al.
+Gender inequalities, exacerbated by AI technologies, illustrate a significant global challenge (Lütz 2023) and exemplify a specific instance of representational harms.
+Gender, a social construct, shapes expectations around masculinity, and femininity, and beyond (Lorber 1996; West and Zimmerman 1987; West and Fenstermaker 1995).
+Gender ideologies adapted to meet various social needs (Martin 2004) and are a major component in structuring individual, interactional, and institutional power relations (Risman 2004).
+A dominant gender construct is sexual objectification, where women are viewed as physical objects and valued for their sexual fulfilment (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011).
+This along with other gender stereotypes affect the performance of real world applications, including online advertising (Sweeney 2013), toxic language detection (Park, Shin, and Fung 2018), machine translation and language technologies (Font and Costa-Jussà 2019; Dev et al.
 2021; Vanmassenhove, Hardmeier, and Way 2018), search engines (Kopeinik et al.
@@
 Stereotypes & Amplification of Societal Biases in T2I Systems
-The representational harm phenomenon in T2I models has been clearly identified (Hao et al.
-2023; Bird, Ungless, and Kasirzadeh 2023; Bianchi et al.
-2023; Cho, Zala, and Bansal 2023; Ungless, Ross, and Lauscher 2023; Wolfe et al.
-2023; Yu et al.
-2022; Jain et al.
-2021; Choi et al.
-2020).
-Existing work often focuses on two areas: (1) identifying the kinds and types of harmful representations or stereotypes across different social attributes and axes of identity, and (2) evaluating and mitigating representational societal harms through measurement and technical intervention.
-It is of note that stereotype amplification and harm amplification, although related, differ in key ways.
-Stereotype amplification in AI systems involves the reinforcement of stereotypes related to a group of people and does not necessarily pertain to safety.
-Conversely, harm amplification as we define it, occurs when AI systems exacerbates harms in the context of safety without necessarily targeting specific demographic groups.
-While stereotype amplification has been well studied, the formalization of harm amplification, to our knowledge, is novel.
-Furthermore, our understanding of harm amplification is related, but distinct, from extant work on “bias amplification” examining how patterns of bias appear in training data, and are amplified statistically in model outputs (e.g., (Zhao et al.
-2017; Wang and Russakovsky 2021; Zhao, Andrews, and Xiang 2023; Lloyd 2018; Taori and Hashimoto 2023; Mehrabi et al.
-2021)).
-In this context, “bias amplification” reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics” (Hall et al.
-2022a, p.
-1), often focusing on binary classification models (e.g., (Leino et al.
-2019; Wang et al.
-2022; Hall et al.
-2022b)) and correlations (Zhao et al.
-2017).
-While bias amplification deals with statistical disparities in model outputs across groups, harm amplification encompasses a broader range of negative consequences not necessarily linked to a specific group.
-Reinforcing Harmful Societal Representations
 Early examination into social stereotype amplification focused on how Generative Adversarial Networks (GANs) exacerbate social biases along axes of gender and skin tone (Jain et al.
 2021; Choi et al.
-2020).
-Jain et al.
-(2021) demonstrate how generated images have a “distributional shift” away from real world data, in which generative AI systems exacerbate extant social inequalities and patterns of over/under-representation.
-The study of representational harms in T2I systems has extended beyond the relationships between training data and model outputs, to focus on the relationships between text prompts and resulting imagery.
+2020) demonstrating ways in which generative AI systems amplify extant social inequalities and patterns of over/under-representation.
 Bianchi and colleagues (2023) delineate and document many examples of the amplification of stereotypes in T2I systems, particularly with respect to racial and cultural stereotypes in unspecified text prompts.
-They show using Stable Diffusion, for example, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p.
-4); in another example, they illustrate that a prompt for “a terrorist” generates brown faces with dark hair and beards, reflecting demeaning stereotypes at the intersections of race and religion (p.
-3).
-Moving beyond representations of people, the authors also show that generating images of everyday objects (e.g., “a house”) reflects Western and middle-class stereotypes (p.
-3).
-Adding specific targeted identity terms (i.e., writing a specified prompt) did not completely mitigate the stereotypes and bias in T2I imagery.
+They show using Stable Diffusion, that T2I systems amplify white racialized notions of beauty, in which “an attractive person” generates faces similar to a “white ideal” (p.
+4) and “a terrorist” generates brown faces with dark hair and beards.
 Qadri et al.
@@
 Bureau of Labor Statistics.
-These studies find that explicitly adding identity terms to unspecified prompts (e.g., “female security guard” (Naik and Nushi 2023)) does not effectively mitigate the presence of stereotypes in generated imagery (Bianchi et al.
-2023).
 Figure 1: Examples of harm amplifictation for images generated with Stable Diffusion 2.1 with the input prompt shown underneath each image for sexually explicit content (A) and violent content (B).
 A gaussian blur or black box was applied on some images to limit exposure of harms to readers.
-Evaluating Representational Societal Biases
-A parallel body of research focuses not just on identifying the presence and kinds of harmful stereotypes that T2I systems can generate, but also evaluating them.
-Luccioni and colleagues (2023) examine whether representational societal harms in each modality amplify or compound one another by evaluating images generated from two synthetic text prompt datasets: one with gender and ethnicity terms and one with profession identifiers.
-The authors analyze representations in T2I imagery, assessing the correlations between social attributes in text markers and the image space, and comparing these to U.S.
-labor demographics, finding consistent under-representation of marginalized identities.
-Wang et al.
-(2023) employ the Text-to-Image Association Test to measure and quantify the presence of societal bias in T2I models while measuring stereotype amplification in occupations across genders.
-Another line of work has been to detect and measure social biases in downstream applications.
-Perera et al.
-(2023) explore social bias with respect to gender, race, and age in the diffusion-based face generation models, finding that diffusion models exacerbate patterns of over- and under-representation, while Saravanan and colleagues (2023) conducts an intersectional analysis of racialized gender stereotypes and professions to support evaluation of T2I models.
-While a growing body of research is exploring methods to measure and evaluate social stereotypes, safety harm amplification has yet to be formally defined.
-Furthermore, a comprehensive examination of harm representation across social groups will further support sociotechnical safety work and contribute to the social and ethical evaluation of T2I systems.
+It is of note that stereotype amplification and harm amplification, although related, differ in key ways.
+Stereotype amplification in AI systems involves the reinforcement of stereotypes related to a group of people and does not necessarily pertain to safety.
+Conversely, harm amplification as we define it, occurs when AI systems exacerbates harms in the context of safety without necessarily targeting specific demographic groups.
+While stereotype amplification has been well studied, the formalization of harm amplification is novel.
+Furthermore, our understanding of harm amplification is related, but distinct, from extant work on bias amplification examining how patterns of bias appear in training data, and are amplified statistically in model outputs (Zhao et al.
+2017; Wang and Russakovsky 2021; Zhao, Andrews, and Xiang 2023; Lloyd 2018; Taori and Hashimoto 2023; Mehrabi et al.
+2021).
+In this context, bias amplification reflects when a model makes “certain predictions at a higher rate for some groups than expected based on training-data statistics” (Hall et al.
+2022a, p.
+1), often focusing on binary classification models (e.g., (Leino et al.
+2019; Wang et al.
+2022; Hall et al.
+2022b)).
+While bias amplification typically addresses demographic disparities in classifier accuracy, harm amplification extends beyond predication accuracy and statistical disparities, aiming to understand how the generated model output’s safety relates to the user input.
 Defining Harm Amplification
-Our paper focuses on examining unsafe amplifications of harms in a specific set of generative AI: T2I systems.
+Our paper focuses on examining unsafe amplifications of harms in a specific set of generative AI - T2I systems.
 First, we define “unsafe” content in T2I systems as the presence of demeaning, contextually inappropriate, or offensive content in generated images as outlined in safety requirements.
 What could be considered “unsafe” or harmful is highly contextual, and needs to be defined within the broader sociotechnical dynamics in which the system is deployed, including system capabilities, context of use, user audiences, and social or regulatory expectations (Rismani et al.
-2023a; Dobbe 2022; Leveson 2020).
-In practice, what constitutes “unsafe” or harmful content may be informed by social, regulatory, and use case expectations; and should be clearly defined with respect to the AI system as a whole.
+2023; Dobbe 2022; Leveson 2020).
 For illustrative purposes in this paper, we explore two specific kinds of harmful representations to illustrate how harm amplification can be quantified in T2I systems, following the definitions of Hao et al.
@@
   • Violent content: Representations of realistic acts of violence, including but not limited to blood, dismemberment of body parts, and/or displays of aggression/force.
-After harms and safety requirements have been established for a specific T2I system and context, harm amplification can be defined as occurring when the output generated image is more harmful than the input text prompt.
+After harms and safety requirements have been established for a specific T2I system and context, we propose the following definition for harm amplification.
+Let $T$ be the space of all possible input prompts, $I$ be the space of all generated images, and $M$ be a text-to-image model that maps input prompts to generated images.
+Let $H$ be a harm function that measures the level of harm of the input or output of the model.
+For a given input prompt $t\in T$ and its corresponding generated image $i=M(t)$, we define harm amplification as occurring when:
+$H(i)>H(t)+\tau$ (1)
+where
+  • $H(i)$ is the harm level of the output image alone
+  • $H(t)$ is the harm level of the input prompt alone
+  • $\tau$ represents a predefined threshold parameter
+Thus, harm amplification can be defined as occurring when the output generated image is more harmful than the input text prompt.
 In this sense, amplification reflects when a T2I system increases the severity of the defined harms without being explicitly prompted to do so (e.g., sexually objectifying women in generated imagery).
@@
 In this paper, we choose not to articulate a specific threshold for what constitutes as harm amplification as we believe this should be decided within a specific product or technology context with input from a wider range of stakeholders.
-Instead, we offer a methodological framework that demonstrates how to derive thresholds in various deployment scenarios and show in our experimental section how to employ this framework on an example dataset.
+Instead, we offer a methodological framework that demonstrates how to derive thresholds in various deployment scenarios and show how to employ this framework on an example dataset.
 Framework to Identify Harm Amplification
-Quantifying harm amplification is a crucial component in assessing “safety” in T2I systems with respect to defined unsafe content.
+In this section, we introduce three methods to quantify harm amplification. We then evaluate the efficacy of these methods on the human annotated Nibbler evaluation dataset (Quaye et al. 2024).
+Methods
+Quantifying harm amplification is a crucial component in assessing safety in T2I systems with respect to defined unsafe content.
 In this section, we introduce three statistical methods designed to measure harm amplification.
@@
 Method 1: Distribution-Based Thresholds for Discretized Text Harm Scores
-The Distribution-Based Thresholds method utilizes (1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and (2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image).
-Consider a set of text harm scores ($T$) and image harm scores ($I$).
-We first discretize $T$ into $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
-$B_{j}=\{t\in T:l_{j}<t\leq l_{j+1}\},\quad j=0,1,...,n$ (1)
-For each $B_{j}$, the corresponding image harm scores ($I_{j}$) form a distribution ($D_{j}$).
+The Distribution-Based Thresholds method utilizes: 1) a text safety classifier that is applied to the input prompt to produce text harm scores (the harm severity for a given prompt), and 2) an image safety classifier that is applied on the output image to generate image harm scores (the harm severity for a given image).
+Consider a set of text harm scores (${H(t):t\in T}$) and image harm scores (${H(i):i\in I}$).
+We first discretize the text harm scores into $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound of the bucket and $l_{j+1}$ is the upper bound of the bucket.
+$B_{j}=\{t\in T:l_{j}<H(t)\leq l_{j+1}\},\quad j=0,1,...,n-1$ (2)
+For each $B_{j}$, the corresponding image harm scores $H(i):i=M(t),t\in B_{j}$ form a distribution $D_{j}$.
 Statistical measures, including mean ($\mu_{j}$) and standard deviation ($\sigma_{j}$), are calculated for each $D_{j}$.
-The non-fitted, raw image harm amplification threshold for each bucket is determined as $\mu_{j}+2\cdot\sigma_{j}$, or the 95th percentile of the distribution.
-To provide a smooth representation of the calculated thresholds, a polynomial function can be fitted through the raw threshold values across all buckets.
+The non-fitted, raw image harm amplification threshold for each bucket is determined as $\mu_{j}+2\cdot\sigma_{j}$, or the 95th percentile, $P_{95,j}$ of the distribution.
+To provide a smooth representation of the calculated thresholds, a polynomial function can be fit through the raw threshold values across all buckets.
 In our experiments, we fit a first-degree polynomial with the corresponding equation to determine the image harm amplification threshold $HarmAmpThresh_{j}$ for each text harm bucket $B_{j}$:
-$HarmAmpThresh_{j}=\hat{y}_{j}=b_{1}j+b_{0},$ (2)
-$where\quad y_{j}=\mu_{j}+2\cdot\sigma_{j}$
-For example, consider a T2I prompt-image dataset where we have image harm scores $I$ and text harm scores $T$.
-Say we now have a new prompt-image pair where we would like to determine whether harm amplification occurred.
-Given an image harm score (i.e., $i=0.5$) and text harm score (i.e., $t=0.15$), we categorize the text harm score into a specific bucket (i.e., $B_{1}=0.1<t\leq 0.2$).
-For that given bucket, the image harm amplification threshold (i.e., $HarmAmpThresh_{1}$) is obtained by fitting a 1-degree polynomial to the 95th percentile threshold for a prior image harm distribution in each bucket.
-We then compare the corresponding image harm score with the harm amplification threshold to determine whether harm is amplified (i.e., $i>HarmAmpThresh_{1}$).
+$HarmAmpThresh(j)=b_{1}j+b_{0}$ (3)
+where $b_{1}$ and $b_{0}$ are the coefficients of the fitted polynomial, determined by minimizing the least squares error:
+$\min_{b_{1},b_{0}}\sum_{j=0}^{n-1}(P_{95,j}-(b_{1}j+b_{0}))^{2}$ (4)
+Harm amplification is determined to occur for a pair $(t,i)$ if $t\in B_{j}$ and $H(i)>HarmAmpThresh(j)$.
 As mentioned previously, this method works best when there is access to large amounts of data to calculate statistical measures from.
@@
 Utilizing text and image harm safety scores similar to Method 1, the Bucket Flip method provides a more direct comparison of discretized harm categories across image and text.
-Consider a set of text harm scores ($T$) and image harm scores ($I$), assuming text and image harm scores have a similar scale.
-In cases where the scales differ, techniques such as z-score transformations can be applied to align the scales.
-We can then discretize $T$ and $I$ into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 1).
+Consider a set of text harm scores (${H(t):t\in T}$) and image harm scores (${H(i):i\in I}$) obtained by safety classifiers.
+Assuming that the classifiers are well aligned, we can discretize text and image harm scores into the same $n$ even buckets ($B_{0},B_{1},\dots,B_{n}$) where $l_{j}$ is the lower bound and $l_{j+1}$ is the upper bound of the bucket (see Equation 2).
 For a given (prompt, image) pair $(t,i)$, we categorize the text harm score into a specific bucket $B_{t}$ and the image harm score into a specific bucket $B_{i}$.
 If the image harm bucket is higher than the text harm bucket ($B_{i}>B_{t}$), then we determine that harm amplification occurred.
-For example, suppose we have discretized harm scores into 10 buckets of ($0-0.1,0.1-0.2,\text{…},0.9-1.0$).
-Given an image harm score (i.e., $i=0.45$) and text harm score (i.e., $t=0.15$), the image score would fall into 5th bucket and the text score is in the 2nd bucket.
-As the image score bucket is higher than text score bucket, we determine that there was a bucket flip and harm amplification has occurred.
 Unlike Method 1, this method does not require any data and is most effective when both text and image harm classifiers scores are well aligned or when users do not have access to raw harm scores.
-In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical approach for identifying harm amplification.
+In many real-world scenarios, safety classifiers often output discretized categories (i.e., low harm, medium harm, high harm) rather than raw harm scores, making the Bucket Flip method a practical and easy approach for identifying harm amplification.
 Method 3: Image-Text Co-embedding Based Harm Scores
@@
 Large pre-trained image-text models, such as CLIP (Radford et al.
-2021), embed images and texts such that the cosine distance between corresponding text and images is minimized in the embedding space.
-Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s embedding, $i$, is to a harmful concept in the embedding space than the prompt’s embedding, $p$, is.
+2021), embed images and their corresponding text labels such that the cosine distance between corresponding text and images is minimized in the embedding space.
+Taking advantage of the co-embedding space, harm amplification can be measured by how much closer a generated image’s CLIP embedding, $z_{i}$, is to a harmful concept in the embedding space than the prompt’s CLIP embedding, $z_{t}$, is.
 Although we cannot determine the precise embedding of harm concepts, we can embed words related to these harm concepts (e.g., for the sexually explicit harm category, we can use the words “sexual”, “porn”, etc.
-- see Appendix, Table 2), defining their text embeddings as $h_{0},h_{1},h_{2},...h_{k}$.
-These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space.
-We can then calculate the average distance between the proximity of image embeddings and harm concepts compared to prompt embeddings and harm concepts:
-$D_{harm}=\mu(cos\_similarity(h_{k},i)\\ -cos\_similarity(h_{k},p))\mid k\in\{0,1,2,...,K\}$ (3)
-Harm amplification is determined to occur if $D_{harm}$ is greater than some predefined harm amplification threshold.
+- see Appendix, Table 2), defining the texts as $\{h_{k}\}_{k=1}^{K}$ and their subsequent embeddings as $z_{h_{k}}$ These harm word embeddings are meant to approximate where harm concepts such as sexually explicit may lie in the CLIP embedding space.
+We can define our harm function $H$ by calculating the average cosine similarity between a given text or image embedding and harm concepts:
+$H(x)=\frac{1}{K}\sum_{k=1}^{K}cos\_sim(z_{h_{k}},z_{x})$ (5)
+where $x$ can denote a text or an image and $cos\_sim(z_{h_{k}},z_{x})=\frac{\mathbf{z}_{h_{k}}\cdot\mathbf{z}_{x}}{\|\mathbf{z}_{h_{k}}\|\|\mathbf{z}_{x}\|}$.
+Harm amplification is then determined to occur for a pair $(t,i)$ if:
+$H(i)-H(t)>\tau$ (6)
+where $\tau>0$ is a predefined threshold.
 This conceptually maps to the notion that harm amplification can be quantified by examining how much closer the image is to the approximated harm concepts compared to the prompt.
-Experiments
-In this section, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset.
+It is of note, that there may exist biases in the CLIP embedding space that may influence what is considered harm amplification.
+Applying this method using a debiased embedding space (Wang, Zhang, and Sang 2022) should be considered as a future direction.
+Evaluation of Methods
+Here, we demonstrate how to apply our framework empirically using a combination of safety classifiers, embedding models, and measurement dataset.
 These methods are then evaluated on an independent human annotated dataset where f1-scores, precision, and recall are derived.
-Measurement Dataset and Models
-To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories.
-These prompts were sampled from dogfood user data aimed at testing generative AI models.
-By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment.
-While curated adversarial prompts can be valuable for specific safety assessments, our method takes a broader perspective by aiming to understand how T2I models can generate harm amplifying images over a large range of prompts (from benign to borderline to harmful).
-We used Stable Diffusion 2.1 (Rombach et al.
-2022) to generate images for our prompt measurement dataset.
-Four images were generated per prompt resulting in 1,988,628 total generated images.
-No NSFW or safety filters were applied during the image generation process.
-Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
-Safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2.
-Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1.
-For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al.
-2021).
-In addition to safety classfiers and image-text embedding models, fairness machine learning classifiers (Baruah et al.
-2022) that identified probability estimates of gender presentation in images were also used with estimates for each gender (male, female, or undefined) ranging from 0 to 1 in the evaluation data.
 Harm Amplification Measurement Method Sexually Explicit Violence
@@
 Precision, recall, and F1-scores for sexually explicit content (left 3 columns) and violence (right 3 columns) were reported for each of the three measurement methods.
+Method 1 Measurement Dataset
+To derive statistics for Method 1, we obtained a measurement dataset consisting of 497,157 prompts representing various demographics and prompt categories.
+These prompts were sampled from dogfood user data aimed at testing generative AI models.
+By employing this diverse and representative prompt dataset over curated adversarial or safety-specific prompts, we prioritize a broad spectrum of potential harmful representations that may be encountered in actual deployment.
+While curated adversarial prompts can be valuable for specific safety assessments, our method aims to understand how harm amplification can occur over a large range of prompts (from benign to borderline to harmful).
+We used Stable Diffusion 2.1 (Rombach et al.
+2022) to generate images for our prompt measurement dataset.
+Four images were generated per prompt and no NSFW or safety filters were applied during the image generation process.
+Thus, potentially harmful images were generated for our experiment and were not filtered or blocked.
+Safety Classifiers and Image-Text Models
+Pretrained safety machine learning classifiers were used to provide harm scores for sexually explicit content and violent content for Methods 1 and 2.
+Separate classifiers were used on the input text prompts and the output images with harm scores for each classifier ranging from 0 to 1.
+For the pre-trained image-text model used in Method 3, we chose CLIP Resnet 101 (Radford et al.
+2021).
 Evaluation Data & Metrics
@@
 To derive a confidence score for sexually explicit and violence, we converted the ratings of the annotators into a proportional measure (i.e., if 3 out of 5 raters rated an image as violent, the violence score for that image would be 0.6 or 60%).
-The ground truth labels for harm amplification for each harm were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm).
-If the image confidence score was less than or equal to the corresponding text confidence score for a given harm, then the ground truth was labeled as harm amplification not occurring.
+The ground truth labels for harm amplification were then defined as whether the image confidence score for a given harm was greater than the corresponding text confidence score (i.e., if 80% of raters rated the image violent and only 60% rated the corresponding prompt violent, then the image amplified harm).
 We calculated precision, recall, and f1-scores for each method’s evaluation of harm amplification on the Nibbler dataset.
 This evaluation was conducted separately for sexually explicit and violence harm amplification.
-Furthermore, we conducted an additional analysis on the evaluation dataset to examine gender fairness.
-Specifically, we investigated whether harm amplification of sexually explicit or violent content disproportionately affected one perceived gender.
-Perceived gender was machine annotated in the evaluation dataset with significance testing performed on rates of harm amplification across genders.
-Experimental Results
-Evaluation of Methods in Detecting Harm Amp
+Evaluation of Methods in Detecting Harm Amplication
 To quantify harm amplification using Method 1, Distribution-Based Threshold Method, we discretized the text harm scores for the measurement prompt data into 5 buckets for each harm type: sexually explicit and violence.
 Subsequently, we then obtained an image harm distribution from the measurement data for each of the 5 buckets corresponding to each harm type (see Appendix, Figure 5A).
-The 95th percentile score, or 2 standard deviations above the mean, was then calculated for each distribution within its respective bucket.
-A first-degree polynomial was fitted across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 3B) for each harm type.
+The 95th percentile score was then calculated for each distribution within its respective bucket with a first-degree polynomial being fit across buckets on the 95th percentile score resulting in new fitted thresholds (Appendix, Figure 5B) for each harm type.
 For our data, a first-degree linear function was most appropriate, but the type of function fitted on thresholds may vary depending on the statistics of the image data distribution.
@@
 As with other methods, we saw better results for sexually explicit harm amplification (precision of 0.585, recall of 0.920) than violence harm amplification (precision 0.322, recall 0.566).
-Gender Disparities in Harm Amplification
+Poorer performance on violence across methods could be attributed to less accurate classification models or noisier human labels.
+Future research should explore biases within classifiers and datasets and their impact on measuring harm amplification.
+Examining Patterns of Harm Amplification
+In the previous section, we showed how to apply different methodologies to measure harm amplification.
+Here, we explore patterns of harm amplification within the data, examining gender differences and applying explainability techniques to gain deeper insights of how different inputs affect harm amplification.
+Gender Disparities
 As described in the Related Works section, the amplification of social stereotypes disproportionately affects marginalized communities (Bianchi et al.
@@
 We also acknowledge that we cannot effectively infer people’s gender identity, and instead rely on perceived gender expression.
-Machine annotations of perceived gender (Baruah et al.
-2022) were collected for each image in the Nibbler evaluation dataset.
+Annotations of perceived gender were collected for each image in the Nibbler evaluation dataset using gender classifiers (Baruah et al.
+2022).
 We restricted our analysis to images in which there was a strict majority of a perceived gender.
@@
 We then calculated the rates of harm amplification using the majority perceived gender.
-Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2), in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011).
-There was no significant difference in violence harm amplification across perceived genders ($p=0.492$).
-We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
+Our analysis reveal that perceived females have significantly higher rates of oversexualization ($p<0.001$, see Figure 2),
 Figure 2: Difference in harm amplification rates across perceived genders in the Nibbler evaluation dataset.
@@
 $*p<0.05,\ **p<0.01,\ ***p<0.001$
-Limitations
-While this framework is shown to be effective in identifying harm amplification, we acknowledge the limitations of our methodologies and identify potential areas of future research.
-Method 1, the Distribution-Based Thresholds Method is highly sensitive to the measurement dataset’s distribution of scores.
-If the distribution is skewed (i.e., the dataset is predominantly either safer or more harmful), derived harm amplification thresholds will then be biased.
-Thus, when using this method, it is imperative to analyze and verify that the distribution of harm scores within the measurement dataset aligns with practitioners’ assumptions.
-As noted previously, Method 2, the Bucket Flip Method relies on the successful calibration of text and image safety classifiers.
-Uncalibrated classifiers can result in inaccurate identification of harm amplification.
-Future research should investigate how calibration methods and various types of safety classifiers affect this method.
-Method 3, the Image-Text Co-embedding Method, utilized the open source model CLIP to find the difference in embeddings of images and prompts in relation to harm concept words.
-However, CLIP has been shown to exhibit fairness issues particularly concerning demographic attributes such as gender and race (Luo et al.
-2024; Wang, Zhang, and Sang 2022).
-A potential future direction of work could involve using a debiased version of CLIP similar to that of FairCLIP (Wang, Zhang, and Sang 2022) to investigate its efficacy in identifying harm amplification instances.
-Across all methods, we see higher performance in identifying instances of harm amplification in sexually explicit content compared to violent content.
-Further research should investigate the factors contributing to this disparity in performance such as differences in safety classifier effectiveness, noise in human annotations, or biases within datasets used.
+in line with previous literature positing theories of gendered sexual objectification stereotypes (Fredrickson and Roberts 1997; Szymanski, Moffitt, and Carr 2011).
+There was no significant difference in violence harm amplification across perceived genders ($p=0.492$).
+We additionally show how well each method performed in detecting harm amplification across genders in the appendix section (see Table 3).
+Applying Explainability Techniques
+To gain deeper insights into the mechanisms of harm amplification, we employed explainability techniques such as attribution maps and counterfactual analysis.
+We utilized Diffusion Attentive Attribution Maps (DAAM) (Tang et al.
+2023) to visualize the relationship between input words and output image pixels, leveraging cross attention maps in diffusion models.
+Figure 3 shows attribution maps for two examples in our dataset that exhibited notable harm amplification.
+Figure 3: Attribution maps (Tang et al. 2023) for notable harm amplification prompts.
+For the prompt “Person venting frustration.”, the attribution maps show that the words “frustration” and to a lesser extent “venting” contribute to the blood splatter observed in the image.
+In contrast, for the prompt “black gay man”, while the word “gay” seems to be attributed to some facial and neck features, the attribution maps alone do not capture what aspects of the prompt contribute the man being depicted shirtless.
+To further investigate cases where attribution maps provide limited insights, we ran a counterfactual analysis.
+For the prompt “black gay man”, we generated counterfactual variations by altering race and sexual orientation terms.
+Using the same generation seed, we produced images from these counterfactual prompts (Figure 4).
+Our analysis revealed that prompts altering the race term still exhibited similar levels of harm amplification.
+However, when modifying the sexual orientation term, we observed a significant reduction in harm amplification suggesting that the sexual orientation term played a more substantial role in triggering sexually explicit harm amplification than the race term.
+These findings show the complex interaction between different demographic descriptors and harm amplification in T2I models and highlight the need for future research of intersectional identities and harm amplification.
+Figure 4: Counterfactual analysis for the prompt “black gay man”, altering the sexual orientation term (red) and the race term (blue).
 Discussion
@@
 Using two specific kinds of harmful representations (sexually explicit content and violent content), we provided a framework of different methods to identify instances of harm amplification under different resource and deployment scenarios.
-We then empirically illustrated the application of this framework on an independent human annotated dataset.
+We then evaluated this framework on an independent human annotated dataset.
 Reduced Reliance on Human Annotation
@@
 Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive.
-Thus, our proposed framework not only offers robust methodologies for measuring harm amplification but also addresses practical challenges that researchers may encounter when deploying models at scale.
-Of course, machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al.
+It is of note that machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al.
 2021) — can also reflect social biases.
@@
 Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
-Gender and Harm Amplification
-In analyzing rates of harm amplification in the evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content.
+Understanding Demographic Disparities in Harm Amplification
+In analyzing rates of harm amplification in the human annotated evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content.
 Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images.
 These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender.
-Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts in different sociodemographic groups.
-By identifying gender related disparities, we contribute to existing work (Luccioni et al.
-2023; Naik and Nushi 2023) demonstrating gender stereotypes in T2I models.
-The durability of gender stereotypes in AI systems (Buolamwini and Gebru 2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes.
-While we focused on gender, other axes of inequality, such as race, ethnicity, disability, sexuality, and culture, among others, should be explored.
+Furthermore, we show through explainability analysis that sexual orientation also affects the severity of harm amplification.
+Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts on different sociodemographic groups.
 Future Work
@@
 This work will be pivotal to understanding the root cause of harm amplification, thereby guiding the development of solutions to reduce its impact in T2I models.
-Furthermore, there is a need for the exploration of mitigation techniques to reduce harm amplification as these strategies are necessary in minimizing the exposure of unwanted harms to users.
+Furthermore, there is a need for the expansion of T2I safety mitigation techniques, similar to Safe Latent Diffusion (Schramowski et al.
+2023), to reduce harm amplification as these strategies are necessary in minimizing the exposure of unwanted harms to users.
 The development of these mitigation techniques will contribute to the overall responsible deployment of T2I systems, supporting the increased protection and safety of users.
+Acknowledgements
+We thank Kathy Meier-Hellstern, Alexander Vorontsov, Vinodkumar Prabhakaran, and Atoosa Kasirzadeh for their feedback and suggestions.
+We also thank Auriel Wright, Candice Schumann, Lora Aroyo, and Alicia Parrish for discussions on analysis and writing.
 Appendix A Appendix